A Bibliography of Publications on Computer-Based Spelling Error Detection and Correction

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: https://www.math.utah.edu/~beebe/

17 May 2024 Version 2.57

Title word cross-reference

n [Ull77, Wag74].

-gram [Ull77].

/usr/group [USE82].

12th [Ano92a]. 14th [Pre91]. 15th [Ano92b, Wil94]. 16th [Hes90]. 17th [RW92]. 17th-century [RW92]. 1880-1920 [Car93]. 1981 [ACM81, BS^{+}81, LT81]. 1982 [Mor82, USE82]. 1983 [ACM83, Law85]. 1984 [Boy85, O’S84]. 1985 [ACM85]. 1986 [KW87]. 1987 [IEE87b]. 1988 [ACM88]. 1989 [Ano89]. 1990 [ACM90, Hes90, IEE90]. 1992 [IEE92a]. 1993 [IEE93c]. 1994 [RAR94, Wil94]. 1995 [IEE95]. 19th [RAR94].

21st [RAR94]. 2D [Lav93]. 2nd

[Ano93e, IEE90].

3rd [Ano94c, LT81].

6th [Ano93b].

80 [Hus81]. ’81 [BS^{+}81]. ’87 [IEE87b].

’91 [ASP91]. ’92 [BIP92, IEE92a]. ’93 [Ano93b, Ano93c, IEE93a]. ’93-Nagoya [IEE93a]. ’94 [BJ94, Tro94]. 9th [Ano93a].

= [IEE87b].

Abbreviations [EP82]. accept [IN94]. access [Mil95, WG10]. Accuracy [LMPK92, Sen93]. achievement [AB84]. Achieving [MM92, TWT84]. ACM [ACM81, ACM83, ACM85, ACM88, ACM89, ACM90, GMJ90, BIP92]. Acoustics [IEE95]. Acquired [CMVR87]. Actes [ACM88]. Actes/Proceedings [ACM88]. Actuated [Can81]. Adaptation [Bey88]. Adaptive [RSM07]. Add [MM92]. Add-A-Word [MM92]. Advanced [Ano93c]. Advances [O’S84]. Advantages [DRG11]. Advertisement [Ano85b]. affected [BH92]. affix [BH92]. agents [Fou94]. Agglutinative [Ofl92, OG94]. AI [Mac87, Rob86]. Aid [GMW75, NH84, Pai81, Rob86, KW81]. aided [HRLJ93, Klixx]. Aids [Che81, TL82]. Airline [Dav62]. Airport [IEE90]. Algorithm [AAA19, DC92, HI20, MW94, VB12, Wie86, KW81]. Algorithms [AFCB94, DC94]. almost [Tay86]. alphabet [LC03]. Alphabetic [LO95]. alternative [Ano91]. Alternatives [Car93]. American [Ano84, Hod87]. Analysis [AAAA93, Chi84, IEE93b, SPR90, Sid79, BH92, KC94]. analytic [BCK85]. Animation [PM17]. Annual [ACM85, ACM90, BIP92, BJ94, Pre91, ACM89]. Anticipation [MPG^{+}92]. any [KPB87a, KPB87b]. APL [Ive91]. Appendix [MC75a]. Application [BG84, IK83, Ofl92, IYME93, MW94, SC96, TIAY90]. Applications [AT&T86, LK93, CR95, MV93, Van83, RAR94]. Applied [Ano94a, SW93]. Approach [Bic87, DC94, LCDM15, LM84, Van83]. approximate [DC94, MW94, ZD95]. April [Ano93b, IEE93c]. Arabic [AMS20, AFCB94, AAA19, BH92, ESH89, PM13, RZR^{+}11]. architectures [DGK^{+}90]. Arizona [CBC84]. array [Fid97, Lav93]. Articles [NCNS82]. Artificial [Ano94c, CC89, IEE90, Ano93e, IYME93, O’S84, RAR94, Ano92a]. ASIC [IEE92a].

Aslib [Law85]. ASMC [Ano93c]. Aspell [AH17]. Assistance [FZ88]. Assisted [BJ94, GY68, TWT84]. Association [Ano84, Ano94a, Law85, Par83, USE82]. associative [CVB90, TST93]. AT&T [AT&T86]. Atlanta [ASP91]. Atlantic [GMJ90]. attitude [AB84]. Augmented [Hus81]. August [Ano92b]. Authoring [Hur87]. authorship [Smi83]. Auto [CRD16, LHF^{+}20]. Auto-Completion [CRD16, LHF^{+}20]. AutoCAD [Yar95]. automata [LK93]. Automated [Hei91, EP82, HRLJ93, LHR^{+}92]. Automatic [AFW83, Bic87, Dol86, GR78, LM84, PZ82, PZ84a, Zam80, Cha94, LM87, Ull77, Wot93]. Automatically [Kuk92b]. Automation [BDT94, Boy85]. Avignon [Ano92a]. Aviv [BS^{+}81].

B [MC75a]. Back [Kuk88].

Back-propagation [Kuk88]. backtracking [KC94]. Baidu [LHF^{+}20]. Base [Zam80]. Based [AAAA93, DS94, DC85, Ebi92, JJJT19, LC16, MDM91, YCS15, Cav94, CVB90, DBT^{+}12, WWSG94, PTKJ07, PT07, TV94, Tur81, Wot93]. Baseforms [LM84]. Bases [Bou77, Gon83, WMRB77]. Basic [Thi82]. Basque [CC89, CA89]. behavior [WG10]. Benchmarking [AH17]. best [JVFM95]. Better [AAA19, SW93]. Between [Chi84]. Beyond [Dun81, Mil81, Pet81]. Bibliographic [Bou77]. Bibliography [Pol82]. binary [Ull77]. Biomedical [IEE87b]. biomedicales [IEE87b]. Blind [NH84]. board [CT92]. Boston [Ano93c, USE82]. brain [DBT^{+}12]. brain-computer [DBT^{+}12]. Bratislava [SD94]. Brief [SF15]. Bucharest [Ano93a]. Bucket [CS82]. Buffer [CMVR87]. Bug [Rob86]. Builder [Fly88]. Built [LC24]. Bulgaria [Boy85]. byte [IEE87a].

CA [ACM83, ACM90, VP94]. CALICO [BJ94]. California [CB87, IEE88, IEE93c]. Canada [ACM85, IEE87a, Pre92]. Cannes [ACM88]. Capabilities [Ber86b]. Carolina [BS83]. Case [CMVR87, FW82, SKS96]. catalogs [WG10]. Cedar [Tei84]. Center [ACM89, IEE92b]. Century [Edm82, RW92]. Changes [LMPK92]. channel [Dol88a, Dol88b]. Character [TST93, TV94]. Characterization [PZ83]. Characters [KPB87b, KPB87a, LAS93]. Check [Ano85a, Ano86a, Ano95, CWC15, LCDM15, ZCX^{+}17, MM93]. Checker [AKGH93, Ben85, Cun94, Fly88, KKC04, Ric98, Wie86, YCS15, Alt90, Ano92d, Arc83, Bom82, CEU94, CEU93, Cri88, Rob83, Sil85]. Checkers [AH17, Mit87a, MM90, Ofl92, Ros82, Che87, CVB90, Fur90, Mit87b, HP88]. Checking [Ano93d, Hug04, LLWL15, Tur81, CW04, GDEJ05, LNL99, MW94, NSV^{+}10]. Checks [WS94]. Children [NF92]. Chinese [Cha94, CWC15, HBHC15, WWSG94, LNL99, LLWL15, LCDM15, LC16, YCS15, ZCX^{+}17]. chip [Lav92]. City [GMJ90, IEE93b, Pre91]. class [Van83]. Classic [RB05]. Classification [Pai81]. Clause [Ber86b]. Cleansing [LC24]. clerical [PM87]. Cn [Mea87]. CNIT [IEE92a]. co [Law85]. co-sponsorship [Law85]. COBOL [VD84, LD76]. Code [MF82a, Hug04]. Codes [SGM11]. Cognitive [PMC85]. COLING [Ano92b]. COLING-92 [Ano92b]. Collection [CC89, NCNS82, PZ83]. collections [SF15]. Collegiate [Pet82]. Colorado [SK83]. Command [Ebi92, Ric98, IYME93]. Commercial [CR95]. Commission [Law85]. Common [MC75a]. Commonwealth [ACM89]. Communication [GT84, IEE87a, SW93]. Communications [SF15, Wil82]. Communities [Law85]. Comparison [MM92]. Completion [CRD16, LHF^{+}20, TST93]. Complex [HSSS78]. Components [BKP86]. Compositors [Har91]. Compound [FZ88, Ros82]. Comprehensive [WWD^{+}21]. compressed [Mil95]. compression [Coo05]. Compu [MM92]. computable [EH88]. Computation [MV93, VMA95]. Computational [Ams86, Ano94a, Ano92b]. Computer [ACM89, Ada84, Ano86a, BS82, BRSS94, BJ94, Che81, Chi84, GY68, GHR67, Hus81, Klixx, Law85, Mor82, MC74, MC75b, Pai81, Pet80a, Pet80b, Pol82, SM80, Ter81, Wil82, Ano92d, Ano93a, AB84, CT92, CHJ94, DBT^{+}12, Dam64, Gor85, HRLJ93, LM87, PM87, TWT84, TG83, Tur81, BRSS94, FW85]. Computer-aided [Klixx, HRLJ93]. Computer-Assisted [GY68, TWT84]. computer-based [Tur81]. computer-managed [TG83]. Computer-Tutor [Ter81]. Computerized [Hei72]. Computers [ACM83, BS83, Dev94, LT81, Rad83, LT81]. computing [Pfa86]. Computr [Mea87]. Concurrent [Ano86a]. Conference [ACM83, ACM85, ACM89, Ano93b, Ano93c, Ano94a, Ano94b, Ano94c, BS^{+}81, BRSS94, BIP92, BS83, CC89, CBC84, CB87, GMJ90, Har94b, Hes90, IEE86, IEE87b, IEE88, IEE90, IEE92b, IEE93a, IEE93b, IEE93c, IEE95, INN90, KW87, Law85, LT81, Mor82, O’S84, Sch82, SD94, USE82, VP94, Wil82, Ano92a, Ano92b, Ano93a, IEE87a, Pre91, Pre92, RAR94, ACM90]. conferences [IEE87b]. Confusion [LC16]. Congres [INN90]. Congress [CC89]. conjugation [Bos94]. connaissance [EBR87]. connectionistic [DGK^{+}90]. Consistency [LMPK92]. Consortium [BJ94]. constraints [Sen93]. Construct [Pet82]. Consultants [Chi84]. consumers [Pre92]. containing [OL95]. Context [Büy20, Jac83, MDM91, EF13, Ram94]. Context-Dependent [Büy20]. context-sensitive [EF13]. Continuous [LO95]. continuously [JVFM95]. Contributed [CC89]. Control [Ano93a]. controlled [DBT^{+}12]. Convention [ACM89, IEE92b]. Conventional [CVB90]. Conversion [LO95, Bes94]. cooperative [CW90]. Copenhagen [BIP92]. corpus [WWSG94]. correct [PTKJ07]. Correcting [EBR87, HBHC15, Kuk92b, MF82a, Pet80a, Ram94, Bla60]. Correction [AAAA93, AMS20, AFW83, Ano92c, AAA19, BG84, BA88, Bic87, Büy20, DS94, DC85, Dol86, DC92, Dun81, DLS83, Ebi92, GY68, GHU84, GHR67, JJJT19, LC24, MDM91, Mil81, Mor69, OG94, Pet80b, Pet81, PZ82, Pol82, PZ84b, RZR^{+}11, Sid79, TKS86, Urq84, Wag74, WF74, WWD^{+}21, Yer86, YEMR91, Zam80, Adr95, Ano91, BSY00, BH92, Ber87, Cha94, CG91, Dam64, EF13, EBR87, Fid97, Hei72, IY85, KO84, Kuk92a, Lav92, Lav93, LM87, Mor70, PZ84a, RSM07, RW92, RS81, SC96, SS84, TIAY90, TTK86, TV94, Ull77, VTTC94, VTTC95]. Corrector [AKGH93, IYME93, YF83a]. Correctors [Mit87a, Mit87b]. Corrects [AHZ81]. Correspondences [HHHR66]. cost [TG83]. Could [Edm82]. courseware [DR94]. crawled [RSM07]. Croatian [DRG11]. Cues [HHHR66]. CUG248 [Art89]. CUG360 [McC92]. CUG374 [Law92].

Data [AH17, Bou77, GMJ90, Gon83, Gri83, HSSS78, WMRB77, WL84, Yan87, Zam80]. Database [CFGJ91, CRF92, Hei91, RW92]. Databases [Ber86b, PZ82, Sch82, Jam84]. datasets [Mil95]. deaf [Kuk92a]. December [ACM83]. Decoding [HBHC15, Dol88a, Dol88b]. decryption [Che87]. Deep [AMS20]. deletion [Ull77]. Denmark [BIP92]. Dependent [Büy20, RSM07]. Depth [Boy85]. Design [FW82, PZ84b, Rad83, Yul94, Gor85, TG83]. Designing [WS84, DC94]. Detecting [EBR87, MF82a, Pet80a, WMRB77]. Detection [AMS20, AAA19, MC74, MC75b, PZ82, Pol82, PZ84b, Zam80, Dam64, Hei91, Hei72, Smi83, EBR87]. Determination [LM84]. Deutsche [BN85]. Developing [HSSS78, LC03]. Development [ACM85, BIP92, McI82, McI95, BS^{+}81, TG83]. Device [Gri83]. Diagnosis [LD76, Bos94]. dialect [Sla87]. Dialectal [RZR^{+}11]. Dictionaries [Ano86c, Yan87, RSM07]. Dictionary [AHZ81, Ano85a, Arn83, Bre93, CS82, Nix81, Par83, Pet82, Rad83, RZR^{+}11, SM80, CW04, Yan91]. DIDAC [DR94]. Diego [ACM83, CB87, IEE88]. Digital [GHR67, Gri83]. d’ingénieire [EBR87]. Directions [BS82]. Disambiguation [TV94]. disk [Ano92d, Arn83]. Display [KW87, Thi82]. distance [EH88, MV93, SKS96]. distances [VMA95]. Districts [Car93]. diverse [SHC83]. Diversifying [CRD16]. do [Dit87]. docile [Fou94]. Document [IEE93b, Jam84, KL82, LAS93, NCNS82, VP94]. documents [Ber87, FQA89]. Does [GDEJ05, JS91]. domain [RSM07, SKS96]. domain-dependent [RSM07]. Domains [SPR90]. d’orthographe [EBR87]. Dulles [IEE90]. during [FS92]. Dutch [Bos94]. Dynamic [Ano85a, JVFM95]. Dysgraphia [CMVR87]. Dyslexia [PMC85]. Dyslexic [NF92, LBW92].

EACL [Ano93b]. East [CB87]. Easy

[Rob86]. ECAI [O’S84]. ECAI-84 [O’S84]. Ecrire [GT84]. ed [LC24]. Edit [Ano86a, MV93, SKS96, VMA95]. editing [FK90, FQA89]. editor [HK95, Dun81, Mil81, Pet81]. Editors [Thi82]. Education [BRSS94, CBC84, LT81]. Effect [LMPK92, AB84, WG10]. effectiveness [BSY00]. Efficient [Gon83, JJJT19, Nix81, Wie86, Yan87]. Eighth [ACM85, IEE86]. Electric [Rob83].

Electronic [Bre93, GMW75, Hur87, TL82, Ber87]. Electrotechnical [BS^{+}81]. Electrotechnology [BS^{+}81]. Embedded [Hes90]. Empirical [VD84, PM87]. Engineering [EBR87, Thi82, IEE87a, RAR94]. engineers [Wei82]. English [Adr95, CS82, Coo05, EBR87, Fur90, HR92, Hua87, Klixx, MM93, MC75a, PM87, RW92, RL78, Sko82, Ven76]. Entry [Gri83, SS14, WCD83]. Enumeration [Car93]. Environment [MF82b]. Environmental [SD94]. equipment [Mit85]. equivalents [PTKJ07]. erreurs [EBR87]. Error [AMS20, Bos94, DC85, LD76, Pai81, PZ82, Pol82, WWD^{+}21, Cha94, EF13, LM87, SC96, TTK86, YF83a]. Error-Proneness [LD76]. Errors [AHZ81, AAA19, Bou77, DC92, EBR87, FS92, HBHC15, LD76, Mac87, MF82a, MC74, MC75b, Pet80a, PZ83, PZ84b, WMRB77, Zam80, Ber87, Bla60, Dam64, Dol88a, Hei91, Hei72, LM87, MC75a, OL95, Pet86, Ram94, SS84, Tur81, Ull77, YF83b]. Essential [SM80, Cri88]. Euro [IEE92a]. European [Law85, O’S84, SD94]. Evaluation [TL82, BSY00, HK95, PM87, TG83]. Evidence [CMVR87]. Example [EBR87, DRG11]. exemple [EBR87]. Experience [AH17, Nix81]. Experiment [Hua87]. Experimental [GY68]. Expert [BA88, Ano92a]. Exploration [LMPK92]. Exploring [GF87]. Extending [Ber86b, Ano92d].

Face [Yul94]. Factor [BJ94]. Factory [Ano93c]. familiarity [WG10]. Fast [CS82, DC92, Rob86, SPR90, VMA95, DC94]. Fault [Sch94]. Fault-tolerant [Sch94]. Faulty [MF82b]. features [LHR^{+}92, VD84]. February [ACM89, VP94]. Fifteenth [BIP92]. File [Ano86b, IK83, Yan87, IY85]. Files [Bre93]. Filter [Urq84]. filtering [Cav94]. Finding [ZD95]. Finds [Pro95]. Fingerspelled [PM17]. Fingerspelling [SGM11]. finite [LK93]. First [AHZ81, CB87, Edm82, SD94, BS^{+}81, TWT84]. FITE [PTKJ07]. Flags [Mac87]. Font [KPB87b, KPB87a]. Foolproof [WS94]. foreign [Gor85]. foreign-language [Gor85]. formatter [MM93]. formatting [Jac83]. forming [BDT94]. forms [RW92]. FORTH [Hes90]. FORTRAN [Hei72]. Fourth [Bic87, Ano94a]. Fourth-Generation [Bic87]. Framework [CWC15, Adr95, Ber87, CC89, EF13]. Francaise [GT84]. France [ACM88, Ano92b, IEE86, INN90, Ano92a]. Francisco [ACM90, IEE93c]. French [Bey88, DR94]. Frequency [Bou77, PTKJ07]. Frequency-based [PTKJ07]. Fully [Ano86b]. Function [GHU84, Urq84]. Further [LMPK92]. Future [Ano93c, IEE87a, Yar95]. Fuzzy [IEE93c, IYME93]. Fuzzyness [Sch75].

Geller [Ang90]. Genealogy [Car93].

generalized [OL95]. Generalizing [SKS96]. Generated [Chi84]. Generating [LC16]. Generation [Ano86b, Bic87, Yan91, Kuk88]. Georgia [ASP91]. German [KW81]. Germany [Ano94a]. GIS [ASP91]. GIS-LIS [ASP91]. Glove [Gri83]. Google [LC24]. grade [AB84]. graders [TWT84]. gram [Cav94, Ull77]. grammar [Adr95, HP88, Klixx]. grammars [JVFM95]. Grammatical [EF13, WWD^{+}21]. Grapheme [HHHR66, Bes94]. grapheme-to-phoneme [Bes94]. Graphemic [CMVR87, LMPK92]. graphical [HR92]. Greek [VTTC94, VTTC95]. Group [Law85, USE82, USE82]. guide [Pfa86]. guided [CW90]. Guild [Law85].

hand [HR92]. handbook [Pfa86, Stu82]. Handicapped [Ter81]. handwritten [PM13, SKS96]. Harbor [CB87, IEE88]. Hart [Har91]. Hash [CS82, MF82a, Wie86]. Hash-Bucket [CS82]. Hashing [Flo87, Mil95]. Having [Ano86a]. Head [MT74]. Hebrew [Sch82]. held [CC89, RAR94]. Herndon [IEE90]. Heuristic [Coo05]. Heuristical [Bes94]. Hidden [DS94]. Hierarchical [IK83, IY85]. High [Flo87, Lav92, TTK86, TKS86]. High-Performance [Flo87]. Hindi [JJJT19, SS14, SS84]. historical [RW92]. History [Ven76, FW82]. HMM [PT07]. HMM-based [PT07]. Honour [BRSS94]. Horn [Ber86b]. Hotel [ACM83, ACM88, IEE87a, IEE90]. Houston [Mor82]. Human [BS82, BJ94, LM87]. Human-Computer [BS82, LM87]. Humanities [BS83]. Hunspell [AH17]. Hyatt [IEE90]. Hybrid [LCDM15, VB12, ZCX^{+}17]. Hypercard [Fly88]. Hyphenation [Pet82].

IBM [Mit85]. Icon [Fly88]. ICSLP [Ano94b]. Identification [JS91, CW04, PTKJ07]. IEEE [IEE90, IEE93c, Ano93c, CB87]. IEEE/SEMI [Ano93c]. IFIP [Boy85, LT81]. IJCNN [IEE92b, IEE93a]. IJCNN-91-Seattle [IEE92b]. imagery [DBT^{+}12]. Impact [Bou77]. Implementation [Mil90, HK95]. Implementations [HI20]. implications [Yan91]. improve [TWT84]. Improvement [HHHR66, LHR^{+}92]. Improving [Fur90, Sch82, Sen93]. In-Depth [Boy85]. Includes [Fly88]. incorporate [SKS96]. Incorporating [Par83]. Independent [CFGJ91, JLMG93]. Index [YCS15]. Indian [SC96]. Inductive [WL84]. inflection [CA89]. Information [ACM85, BIP92, Law85, LM84, SD94, Wil82, HK95, SKS96]. Informational [Wil82]. Inforum [ASP91]. INNC [INN90]. insertion [Ull77]. Insights [AH17].

Instruction [BJ94, Hod87, TWT84]. Instrument [TL82]. Integrated [Ano92c, Lav93]. Integrating [SHC83]. Integration [Dal87]. Intelligence [Ano94c, CC89, IEE90, Ano92a, Ano93e, O’S84, RAR94]. Intelligent [AHZ81, YEMR91, YF83a]. interaction [LM87, SC96, BS82]. Interactive [Rad83, RW73]. Interactively [FQA89]. Interface [Gri83, HSSS78, CW90, DBT^{+}12]. Interfaces [DLS83]. International [ACM85, Ano92a, Ano92b, Ano93a, Ano94b, Ano94c, BRSS94, BIP92, Boy85, BS83, CB87, GMJ90, IEE86, IEE87a, IEE88, IEE90, IEE92b, IEE93a, IEE93b, IEE93c, IEE95, INN90, KW87, Pre92, Sch82, WS84, RAR94]. Introduction [LLWL15, Gor85]. Inverted [YCS15]. investigation [VD84]. Invited [CC89]. IPCC [IEE87a]. ISKO [SD94]. Island [CB87, IEE88]. ispell [Ric98]. Israel [BS^{+}81, Sch82]. Issue [LLWL15]. issues [Gor85]. Istanbul [Ano93e]. Italy [O’S84]. IX [RAR94].

Jake [Ang90]. January [ACM90]. Japan [IEE93b]. Japanese [Bre93, Fur90, HK95]. Jerusalem [Sch82]. Jewel [LC24]. Joint [IEE92b, IEE93a, USE82]. Jointly [Law85]. Jose [VP94]. July [IEE88, IEE92b, INN90, LT81, RAR94, USE82]. June [ACM81, ACM85, Ano92a, Ano93e, BIP92, BS83, CB87, Hes90, IEE92a, Mor82, Pre91, Pre92, Sch82]. Just [Cun94].

Kaleidoscope [CW90]. Kansas [Pre91]. katakana [IN94]. Kaypro [Hic83]. Kentucky [ACM89]. Keyboard [Can81, WCD83]. Keywords [Sch75]. Khmer [Har94a]. Knowledge [Dal87, Ebi92, EBR87, SD94, SHC83]. Knowledge-Based [Ebi92]. KONVENS [Tro94]. Korean [KC94, KKC04].

label [GDEJ05]. Language [Ano86c, Ano94a, Ano94b, BH95, Bic87, BJ94, HSSS78, JJJT19, LO95, PM17, Woo85, Ano92a, BDT94, CR95, GF87, Gor85, RL78, Sla87, dO94]. Languages [ACM90, OG94, Wag74, SC96]. Large [LHF^{+}20, WMRB77, Zam80, LK93, Yan91, ZD95]. Large-Scale [LHF^{+}20]. Lattice [HBHC15]. Lausanne [LT81]. Learning [AMS20, Hua87, JJJT19, TL82, WL84, DR94, dO94]. learns [FW82]. Legal [Ano93d]. Legal-Specific [Ano93d]. Lemmatizing [KW81]. Length [Tur82]. Length-segmented [Tur82]. lessons [Klixx]. Letter [Dun81, Mil81, Pet81, BH95]. Lexical [Ber86b, MPG^{+}92, SC96]. lexicology [Ams86]. lexicon [CD01, VTTC95]. lexicons [ZD95]. life [Ano92d]. Lightning [Yer86]. line [RW73]. Linear [Fid97, EH88]. Linguistics [Ano94a, Ano92b]. LIS [ASP91]. LISP [Sla87]. List [McI82, McI95, Pet82, YCS15, Coo05]. lists [Tur82]. Literacy [CBC84]. literary [Smi83]. Literate [Knu92]. Live [FK90]. logical [Ber87]. LOGO [GF87]. London [Wil82]. look [Cri88]. looks [Yar95]. Lookup [RZR^{+}11]. lost [Ano92d]. Louisville [ACM89]. Low [MM92, TWT84, TG83]. Low-Achieving [MM92]. low-cost [TG83]. LPC [BG84]. LRE [Adr95]. Lumpy [Yan87].

machine [EF13]. Macro [LHR^{+}92]. mai [ACM88]. Making [Joh81]. Malay [BSY00]. man [Arc83]. managed [TG83]. Management [GMJ90, Law85, SD94]. Managers [SM80]. Manipulation [ACM81]. Manual [Ano84]. Manufacturing [Ano93c]. Manuscripts [LC24]. Maps [LHF^{+}20]. March [CBC84, IEE93c]. Market [WS84]. Markov [DS94]. Massachusetts [Ano93c]. Master [Pet82]. match [Tay86]. matches [ZD95]. Matching [IK83, VB12, BSY00, DC94].

Mathematics [Dev94]. Matter [Pro95]. Max [Alt90]. May [Ano89, Ano93a, BS^{+}81, GMJ90, KW87]. Me [Cun94]. Meaning [JS91]. Measure [AFW83]. Measures [Kre92]. Mechanism [YCS15]. Mediterranean [BS^{+}81]. Meeting [Ano89, Wil94]. MELECON [BS^{+}81]. membership [MW94]. memory [CVB90]. memory-based [CVB90]. menu [CW90]. menu-guided [CW90]. Message [Che87]. Method [AAAA93, LC16, Mil90, MF82a, WMRB77, IN94, PT07, TIAY90]. methods [BSY00, Bes94, KO84, RW92, WG10]. metric [EH88]. Micro [Art89, Cri88, LHR^{+}92]. Microcomputers [CBC84]. Microprocessored [AHZ81]. MicroSpell [Law92]. Misspelled [Bic87, Dav62, SF15]. Misspellings [Mit87a, Alb67, Mit87b]. mistakes [Dol88b]. mixed [WG10]. MMSFORTH [Mil90]. MO [Pre91]. Model [DC92, Kre92, LC24, ZCX^{+}17, AB89, LBW92, Ram94]. Models [DS94, BH95]. Modern [VTTC94, VTTC95]. Montech [IEE87b]. Montfleury [ACM88]. Montreal [ACM85, IEE87b]. Morphological [AAAA93, SC96, CA89, ESH89, KC94]. Morphologically [Wot93]. motor [DBT^{+}12]. MT [Adr95]. much [Rob83]. Multi [Ano86c]. Multi-Language [Ano86c]. Multiculturalism [Cun94]. Multiple [Ano86a, Dal87, DC92].

N [Cav94, JVFM95]. N-best [JVFM95]. N-gram-based [Cav94]. Nagoya [IEE93a]. Name [CFGJ91, Mil90, AFCB94]. Names [Bic87, CFGJ91, CRF92, Dav62, Ebi92, VB12, JVFM95]. Naming [LMPK92, LM87, Hei91]. Nantes [Ano92b]. National [Mor82, Wil94, Ano89]. Natural [Ano94a, HSSS78, Tro94, Ano92a, CR95]. Naturlicher [Tro94]. NC [BS83]. need [GDEJ05]. Neighborhood [LMPK92]. net [AB89, LBW92]. Netherlands [Ano93b]. Network [INN90, Ano93e, Hei91, IYME93, Kuk92a, TV94]. network-based [TV94]. Networks [Boy85, CB87, IEE88, IEE92b, IEE93a, JVFM95, Boy85]. Neural [AB89, CB87, IEE88, IEE92b, IEE93a, INN90, LC24, Ano93e, JVFM95, LBW92, TV94]. Neuropsychological [PMC85]. neuropsychology [AB89]. newspaper [KW81]. next [BD95]. nineties [Pre91]. ninth [RAR94]. NJ [GMJ90]. No [Pro95]. Non [CW04, GT84, BCK85]. non-semantic [BCK85]. Non-Verbale [GT84]. Non-word [CW04]. Nonword [LMPK92]. normal [LBW92]. normalized [MV93, VMA95]. North [BS83]. note [Pet86]. Notes [Car93, Pol82, Ven76]. nothing [BD95]. November [ASP91, IEE87b, IEE90, Law85]. novembre [IEE87b]. Number [PM17].

obtained [PTKJ07]. OCR [LC24, TIAY90, TKS86]. OCR-ed [LC24]. October [ASP91, Ano93c, Ano94a, IEE86, IEE87a, IEE93b]. Office [Boy85, GR78, SM80, Mit85]. Offline [PM13]. Omnifont [BKP86]. on-line [RW73]. Online [Ano89, Wil94, DRG11, WG10]. OPACs [Sch94]. operators [EP82]. Optically [Can81]. Optimization [IY85]. Order [Wag74]. Order- [Wag74]. Oregon [ACM81]. Organization [IK83, SD94, IY85]. ORTHO [DR94]. ORTHO-DIDAC [DR94]. Orthographe [GT84]. Orthographic [LMPK92]. orthography [RL78]. Other [Yan87]. Oxford [Har91].

Pacific [Ano94c]. Packed [Lyo78]. Page [BKP86]. Palais [INN90]. paper [WL83]. Papers [ACM90, CC89, KW87, MM93]. Paris [IEE86, IEE92a, INN90]. parser [CA89]. Parsing [Ofl92]. Part [Bre93]. Pascal [KP81]. Passenger [Dav62]. password [MW94]. Pathways [Wil82].

Patient [FS92]. Pattern [IEE86, OL95]. Patterns [DC85]. PC [ACM88, HP88]. PDAs [LC03]. Pearls [Ben85, Ben86]. Pennsylvania [RAR94]. perfect [Mil95]. Performance [Flo87, Lav92]. person [Bom82]. Personal [ACM83, Rad83, VB12, CT92, Ive91, Pfa86]. Personalized [LHF^{+}20]. Personnel [SM80]. Perspectives [CC89]. Phoneme [HHHR66, WL84, Bes94, TTK86]. Phoneme-Grapheme [HHHR66]. phonemes [Yan91]. Phonemic [LM84, Sko82]. Phonetic [HI20, Wot93]. Phonological [PMC85]. pilot [Cha94]. Pin [MT74]. Pisa [O’S84]. Plagued [Rob86]. Plus [CBC84]. POI [LHF^{+}20]. Poor [Mit87a, Arc83, Bom82, Mit87b]. Portland [ACM81]. Possible [BG84, MC75a]. Postal [SGM11]. Pour [GT84]. practical [Pfa86]. practice [SW93]. Prediction [SS14]. Predictive [Ano92c, DBT^{+}12, KC94]. Preparation [NCNS82, KL82]. Presented [ACM90, CC89]. Press [Har91]. PRICAI [Ano94c]. PRICAI-94 [Ano94c]. Principles [ACM90, Thi82]. Printed [KPB87b, KPB87a, LAS93]. Probabilistic [CWC15, KO84]. Probability [CG91]. Problem [WF74, Ram94]. problem-solving [Ram94]. Problems [Yul94, Ano92d]. Proceed [JS91]. Proceedings [Ano93c, Ano94c, BRSS94, BJ94, Boy85, IEE93a, Law85, SD94, USE82, Wil82, Wil94, ASP91, Ano89, Ano94a, BS^{+}81, BIP92, IEE92a, LT81, O’S84, RAR94, ACM81, ACM88, GMJ90, IEE90, IEE93b, IEE95, Sch82, IEE86]. process [BDT94]. processes [AB89, Sch94]. Processing [Ano86c, Ano94a, Ano94b, Ber86b, IEE95, Tro94, Chr87, CR95, MCM86, SC96, Stu82, Woo85, FW85]. Processors [Ano86a]. Prof [BRSS94]. Professional [IEE87a]. proficiency [TWT84]. program [Ams86, Bla60, PM87, RS81, SS84].

Programming [ACM90, Ben85, Ben86, Knu92, Mor69, Sla87]. Programs [Fly88, MM92, NF92, Pet80a, Pet80b, AB84, Hei72, Mor70, SK83, Van83]. Project [Bre93, Adr95]. Prolog [Ber86a]. Proneness [LD76]. pronunciation [HK95, HRLJ93, LHR^{+}92]. pronunciation-information [HK95]. Proof [WL83]. Proof-reading [WL83]. propagation [Kuk88]. Prospect [Hod87]. Prototype [BA88]. Proximity [Mac87]. Proximity-Scan [Mac87]. Pseudohomophone [LMPK92]. Psychological [Ano84]. public [WG10]. Publication [Ano84].

Qualified [Ano86b]. Quebec [ACM85]. Queries [MF82b]. Query [CRD16, LHF^{+}20, CW90]. question [Asi82].

R [HI20]. R13 [Yar95]. Raed [Mea87]. Rajaraman [BRSS94]. Raleigh [BS83]. RAM [Edw87]. RAM-resident [Edw87]. Random [Mil95]. Ranking [LCDM15]. Reaction [Kre92]. Reactive [WCD83]. Reader [BKP86, LMPK92, BCK85]. Readers [Har91, CHJ94]. reading [Van87, WL83, PMC85]. Readings [AT&T86]. Real [AAA19, LO95, JVFM95, Ram94]. Real-time [LO95, JVFM95]. Real-Word [AAA19, Ram94]. Reclaiming [Ano92d]. Recognition [Dal87, IEE86, IEE93b, KPB87b, SGM11, JLMG93, JVFM95, KPB87a, LAS93, OL95, PM13, PT07, SKS96, Sen93, SHC83, TV94, TST93, VP94, Yan91]. recognizer [BH95]. Recognizing [DGK^{+}90]. Record [ACM90, Dav62, IEE87a]. Reduction [WL84]. reflexive [dO94]. Regular [Wag74]. Reject [Ano95]. Remedial [Pai81]. Remediation [NF92]. Representation [GHR67, Sko82]. representing [LK93]. Rescoring [YCS15].

Research [BIP92, LNL99, Yul94, Ams86, ACM85]. resident [Edw87]. RESORT [Kra84]. Responsiveness [Sch82]. results [SW93]. REtrieval [Har94b, ACM85, BIP92, CFGJ91, Dav62, MF82b, BSY00, WG10]. Retrospect [Hod87]. Revelation [Mac87]. reversal [Ull77]. review [Klixx]. revised [RL78]. right [Sil85]. Rim [Ano94c]. road [Pou84]. robust [BH92]. Rochester [Hes90, Hes90]. Role [CMVR87]. Romanian [CEU93, BDT94, CEU94]. ROSE [Van87]. ROWS [Van87]. Rules [Har91, Hua87, WL84, EP82, PTKJ07, YF83b, Yan91]. Russell [Mil90]. Russell-Soundex [Mil90].

San [ACM83, ACM90, CC89, CB87, IEE88, IEE93c, VP94]. Scale [LHF^{+}20]. Scan [Mac87]. Scatter [Lyo78]. scholar [Pfa86]. Scholarly [PZ83, PZ84b, PZ84a]. Science [ACM89, IEE93b, Ano93a, MM93]. Scientific [KW87, PZ83, PZ84b, PZ84a]. scientists [Wei82]. scoring [CG91]. screen [WL83]. Scripting [RB05]. Scripts [Chi84]. Search [CS82, LHF^{+}20, Sch94, WG10]. Searches [Sch75]. Searching [CS82, Flo87, Fou94, Gon83, RW92, SF15]. Seattle [IEE92b]. Sebastian [CC89]. SECC [Adr95]. Second [CC89, IEE93b, MT74, Sch82, dO94, Har94b, IEE93c]. Secretaries [SM80]. security [MW94]. Seed [LC16]. Seed-Based [LC16]. Segmentation [LAS93, LNL99]. segmented [Tur82]. Selected [KW87]. selectively [JVFM95]. Selfspell [NF92]. semantic [BCK85]. SEMI [Ano93c]. Semiconductor [Ano93c]. sensitive [EF13, Jac83]. Sept [CC89, SD94, Wil82]. September [Boy85, O’S84]. Sequence [Büy20, SGM11, Kuk88]. Sequence-to-Sequence [Büy20]. sequences [TTK86]. Sequential [Dol88a, Dol88b]. Sets [LC16].

Seventeenth [ACM90, ACM89]. Seventh [Pet82]. Shell [RB05]. Sheraton [CB87, IEE87a, IEE88]. Sichere [BN85]. SIGIR [ACM85, BIP92]. SIGMOD [GMJ90]. Sign [LO95, PM17, HR92]. Signal [IEE95]. Signed [HR92]. Signs [PM17]. SIGOA [ACM81]. SIGPLAN [ACM81]. SIGSMALL [ACM88]. SIGSMALL/PC [ACM88]. Silbentrennung [BN85]. Similar [IK83]. Similar-String [IK83]. Similarity [AFW83, Sch75, Sid79, Alb67, BSY00]. Simplified [Adr95]. single [SS84]. Sinhala [PM17]. Sinnentsprechende [BN85]. Sixth [O’S84, Wil82, AB84, BS83]. sixth-grade [AB84]. Size [KPB87b, LMPK92, KPB87a]. Skilled [Kre92]. Skipping [GHU84]. Slovakia [SD94]. Small [ACM83, SPR90, Van83]. smarter [Joh81]. Smoothing [KKC04]. Society [Wil82]. Sofia [Boy85]. Software [BD95, Ber86b, Hic83, KP81, USE82, WS84, GDEJ05, MM93]. Solution [GR78]. solving [Ram94]. Some [Mit85, Pol82]. Sound [JS91, LMPK92, MPG^{+}92, BCK85, Van87]. Soundex [Car93, Mil90, Pro95]. Soundspel [RL78]. source [Hug04]. Sources [Dal87, SHC83]. Space [Nix81, Ano92d]. Spacing [KKC04]. Spain [CC89]. Speak [TL82, FW82]. Speaker [CFGJ91, JLMG93]. Speaker-Independent [CFGJ91, JLMG93]. Special [LLWL15]. Specialized [Yan87]. Specific [Ano93d]. Specification [Ano86b]. Specified [Ano86b, Ano86d]. Speech [CRF92, GMW75, IEE95, LO95, Tro94, FW85, WWSG94, PT07, Woo85, Yan91]. Speed [Edm82, TKS86, TTK86]. Speed-Spelling [Edm82]. Spell [AH17, Ano85a, Ano86a, Ano93d, Ano94d, Ano95, Art89, Bey88, Cun94, LLWL15, Pro95, SPR90, TL82, WS94, Alt90, CW04, Dit87, FW82, GDEJ05, LNL99, NSV^{+}10, HRLJ93, LHR^{+}92, MM92]. spell-checking [GDEJ05]. Spellchecker [DGR09]. spellchecking [DRG11]. SpellCode [Ang90]. Spelled [CRF92, BH95, JVFM95]. Speller [Ano91, CT92, SM80, DBT^{+}12, dO94]. Spellers [Mit87a, Mit87b]. Spelling [AHZ81, AAAA93, AKGH93, AFW83, Ano86d, Ano92c, BG84, BH92, Ben85, Ber86a, BA88, Bou77, Büy20, CMVR87, CWC15, CS82, DS94, DC85, Dol86, DC92, Dun81, DLS83, Ebi92, Edm82, EBR87, Fly88, FS92, FZ88, GHU84, GHR67, GMW75, GR78, HHHR66, Hod87, HI20, HBHC15, Hua87, JJJT19, JS91, KO84, Kre92, Kuk92a, KKC04, LMPK92, LCDM15, LC24, MDM91, McI82, McI95, MM93, Mil81, Mit87b, Mit87a, MPG^{+}92, MF82a, Mor69, Mor70, MM90, NH84, NF92, Ofl92, OG94, Pai81, Par83, Pet80a, Pet80b, Pet81, PZ82, Pol82, PZ83, PZ84b, Rad83, Ric98, Ros82, RZR^{+}11, Sid79, Sil85, Sko82, Urq84, Ven76, WMRB77, Wie86, WL84, YCS15, Yer86, YEMR91, Yul94, Zam80, ZCX^{+}17, Ano91, Ano92d, Arc83, AB84, Asi82, AB89, BSY00, Ber87, Bla60, Bom82]. spelling [Bos94, BCK85, Cha94, Che87, CVB90, CG91, CEU94, CEU93, Cri88, Dam64, DR94, Dol88a, Dol88b, Fid97, Fur90, Hei72, HP88, Hug04, IYME93, IY85, JLMG93, WWSG94, Klixx, Lav92, Lav93, LC03, LBW92, MC75a, PT07, PZ84a, PM87, Ram94, RW92, RS81, Rob83, SC96, SS84, TIAY90, TWT84, TG83, TV94, Tur81, VTTC94, VTTC95, Van87, WG10, YF83a, YF83b, Yan91, dO94]. spelling-checker [Ano92d, CEU94]. spelling-correction [BSY00, RW92]. Spelling-Sound [LMPK92]. Spelling-To-Phoneme [WL84]. spelling-to-sound [BCK85]. spelling-word [WWSG94]. Spellings [CFGJ91, Dal87, Ros82]. Spoken [Ano94b, CRF92]. Sponsored [Law85]. sponsorship [Law85]. Sprache [Tro94, BN85]. squeeze [CD01]. Standard [GT84]. Standardizing [SW93]. State [BS83, CBC84, IEE92b, SK83]. Statistical [MCM86, WMRB77, Bes94, EF13]. statistics [WWSG94]. Status [MPG^{+}92]. Steps [Hic83]. Stochastic [KKC04]. Stockholm [KW87]. Store [Nix81]. Strategic [MPG^{+}92]. Strategy [Kre92, JVFM95]. street [Hei91]. Strictness [SPR90]. String [Ada84, Alb67, IK83, TKS86, WF74, BSY00, DC94]. string-similarity [BSY00]. String-to-String [WF74]. strings [BH92, EH88, OL95]. Structure [Yan87, Sch94]. structured [FQA89]. Structures [Ofl92, RW73]. Students [MM92, AB84, PM87]. Studies [GY68, PMC85]. Study [LD76, Cha94, SKS96, WG10]. Stuttgart [Ano94a]. style [Adr95]. Stylometry [Smi83]. Sublexical [MPG^{+}92]. substitution [Ull77]. Supervised [JJJT19]. Supplement [Ano85a, Ano86d]. Support [Ano86d, dO94]. Surface [PMC85]. Survey [NCNS82, WWD^{+}21, PM13]. Sweden [KW87]. Switzerland [LT81]. Symposium [ACM81, ACM88, ACM90, Ano93e, BJ94, Boy85]. System [AT&T86, Ano86a, Ano92c, BA88, Can81, Dav62, GT84, GR78, Hus81, JJJT19, LO95, PZ84b, SS14, YEMR91, Ano91, Bos94, Chr87, ESH89, HRLJ93, HR92, LHR^{+}92, TIAY90, TG83, TV94, VTTC94, Wei82, WG10]. Systems [BRSS94, Hes90, IEE93c, Mor69, NCNS82, Sid79, Ano92a, Ano93a, EP82, Mor70, Sen93]. systolic [Lav92, Lav93].

T [Sla87]. Tables [Lyo78]. tag [NSV^{+}10]. Take [Ric98]. Tale [MM90]. TALISMAN [BA88]. talk [FW82]. Tasks [Ano85a]. TC [Boy85, LT81]. TC-3 [LT81]. Teaching [Hus81, Pai81, EP82, HRLJ93]. Technical [Law85, MC75a]. Technique [CS82, Dam64, Ull77]. Techniques [Kuk92b]. technologies [IEE87b].

Technology [Mac87, WCD83, Pre91, Pre92]. Tel [BS^{+}81]. Tel-Aviv [BS^{+}81]. telecommunications [Kuk92a]. Telephone [CRF92, JLMG93, JVFM95]. Tempe [CBC84]. terminology [Buc93, SW93]. Terms [LMPK92]. Testing [Kre92, PM87]. Texas [Mor82, TL82]. Text [ACM81, Ano86c, GY68, Gon83, Har94b, Kuk92b, LNL99, PZ83, PZ84b, SS14, WCD83, Cav94, Chr87, FK90, HR92, MCM86, MM93, PM13, PZ84a, RSM07, RW92, SKS96, Sen93, SHC83, Tur81]. Texts [AAA19, MF82b, BSY00, BH92, KW81, WL83]. Textual [PZ82]. Thai [PT07]. their [Dal87, Yan91]. Theme [Ano93c]. Theoretic [LM84]. theory [SW93]. Therapy [GMW75]. thesaurus [RW73]. Three [MM90]. Ths [Mea87]. Tibetan [LC24]. Time [Kre92, EH88, JVFM95, LO95]. tolerant [Sch94]. Tolerating [FS92, Sch75]. tool [CHJ94, Cri88]. Toolkit [Mac87]. Tools [Hur87, IEE90, KP81, USE82, HP88, Law85]. topologies [Kuk88]. Toronto [Pre92]. touching [LAS93]. tour [Tei84]. Trade [IEE92b, Law85]. traditional [OL95]. trained [JVFM95]. transcription [Wot93]. transformation [PTKJ07]. Translating [Law85]. Translation [Law85, MPG^{+}92, BCK85, EF13, HR92, WWSG94, PTKJ07]. Translator [Law85, HR92, Mit85]. transliteration [AFCB94]. transposition [OL95]. TREC [Har94b, Cav94]. TREC-2 [Har94b, Cav94]. Trigram [AFW83]. Tsukuba [IEE93b]. Turbo [Yer86]. Turkey [Ano93e]. Turkish [AKGH93, Ano93e, Büy20, Ofl92]. Tutor [Pro95, Ter81]. Tutoring [YEMR91, Bos94]. Twenty [Edm82]. Twenty-First [Edm82]. Two [Kre92, AB84]. Two-Strategy [Kre92]. Type [Yer86]. Typesetting [Har94a]. typing [Pet86]. Typist [Rob86]. typo [MC75a]. Typographical [MC74, MC75b, Sen93, Tur81]. understanding [WWSG94, Woo85]. undetected [Pet86]. Unfold [GHU84]. Units [KW87]. University [BS83, CBC84, Hes90, Sch82, Har91, SK83]. UNIX [AT&T86, Chr87, KL82, Ebi92]. unknown [Fou94]. Unorthographic [GY68]. Unstructured [Gon83]. USA [IEE90, RAR94]. Usability [Sch82]. Use [GMW75, Pet82, TWT84]. USENIX [USE82]. User [Ano86b, Ano86d, DC85, DLS83, Thi82]. User-Specified [Ano86d]. Users [Chi84, USE82, Fur90]. Uses [Mac87]. Using [AFW83, Car93, GHR67, KKC04, NF92, SK83, BSY00, CFGJ91, EF13, KO84, MM93, Ram94, RW92, Sen93, Van83]. Uspell [McC92]. usr [USE82]. Utilities [Edw87]. Utility [Fly88]. Utilization [VTTC95]. utilizing [HK95]. Utrecht [Ano93b]. UTTAM [JJJT19].

V [BRSS94]. v2.0 [Law92]. VA [IEE90]. Validation [FS92]. value [Mil95]. Variant [Buc93, Ros82]. variants [IN94]. VCATS [Hus81]. Vector [GHR67]. Ventura [Alt90]. Verarbeitung [Tro94]. verb [CA89]. Verbale [GT84]. verbs [Bos94]. Verification [Ber86a, GR78]. Very [Gon83, DC94]. Video [SGM11]. view [Ive91]. Virtual [Wie86]. Visual [Hus81]. vocabularies [LK93]. voice [LC03]. voice-spelling [LC03]. VP [Alt90]. VP-to-the-Max [Alt90]. VR [LO95].

WA [IEE92b]. warning [GDEJ05]. Washington [IEE92b]. Way [Nix81]. Web [RSM07]. Web-crawled [RSM07]. Webster [Pet82, Sil85, MT74, Rob83]. well [Dit87]. Westgate [ACM83]. Whole [BCK85]. Whole-word [BCK85]. Winnipeg [IEE87a]. within [CC89]. without [CW04, KC94]. Word [AAA19, Ber86b, HBHC15, JS91, KKC04, McW82, Ofl92, Par83, SS14, Sid79, BDT94, BCK85, CW04, Coo05, WWSG94, Ram94, RW92, Sil85, Stu82, MM92]. word-forming [BDT94]. word-forms [RW92]. Word-Spacing [KKC04]. Wordprocessing [Ano92c]. Words [Dal87, FZ88, Kuk92b, PM17, Ros82, DGK^{+}90, Fou94, MC75a, SS84, TST93, Ull77]. Wordz [Tay86]. Work [Edm82, KW87, KW87]. Workbench [SK83, Van83]. Workshop [Ano93c]. World [CC89, LT81]. WP [Joh81]. Writer [HP88, SK83, Van83]. Writers [Che81, CHJ94]. Writing [Chi84, Rob86, Wei82].

York [Ano89]. Yur [Mea87].

References

Azmi:2019:RWE

[AAA19] A. M. Azmi, M. N. Almutery, and H. A. Aboalsamh. Real-word errors in Arabic texts: a better algorithm for detection and correction. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 27(8):1308–1320, August 2019. ISSN 2329-9290.

Aduriz:1993:MAB

[AAAA93] I. Aduriz, E. Agirre, I. Alegria, and X. Arregi. A morphological analysis based method for spelling correction. In Anonymous [Ano93b], pages 463–?? ISBN 90-5434-014-2. LCCN ????

Arhar:1984:ETC

[AB84] M. Arhar and C. Beverstock. The effect of two computer spelling programs on the attitude and achievement of sixth-grade students. In Camuse et al. [CBC84], pages 1–6. ISBN 0-88175-077-8. LCCN LB1028.5 .M53 1984.

Auernheimer:1989:NNM

[AB89] Brent Auernheimer and Alison Butler. Neural net model of the neuropsychology of spelling processes. In ACM [ACM89], pages 430–?? ISBN 0-89791-299-3. LCCN QA75.5 .A1371 1989.

ACM:1981:ASS

[ACM81] Proceedings of the ACM SIG-PLAN SIGOA Symposium on Text Manipulation, Portland, Oregon, June 8–10, 1981. ACM Press, New York, NY 10036, USA, 1981. ISBN 0-89791-043-5. LCCN QA76.7 .S54 v.16:6. ACM SIGPLAN Notices, v. 16, no. 6, (June 1981).

ACM:1983:ACP

[ACM83] 1983 ACM Conference on Personal and Small Computers: Westgate Hotel, San Diego, CA, December 7–9, 1983. ACM Press, New York, NY 10036, USA, 1983. ISBN 0-89791-123-7. LCCN QA 76.5 S53 v.6 no.2.

ACM:1985:RDI

[ACM85] Research and Development in Information Retrieval: Eighth Annual International ACM SIGIR Conference, Montreal, Quebec, Canada, June 5–7, 1985. ACM Press, New York, NY 10036, USA, 1985. LCCN Z699.A1 I659 1985. US$23.00.

ACM:1988:APS

[ACM88] Actes/Proceedings, Symposium 1988, ACM SIGSMALL/PC, Hotel Montfleury, Cannes, France, 4–6 mai, 1988. ACM Press, New York, NY 10036, USA, 1988. LCCN QA 76.6 A28 1988.

ACM:1989:SAA

[ACM89] ACM, editor. Seventeenth annual ACM Computer Science Conference, February 21–23, 1989, Commonwealth Convention Center, Louisville, Kentucky. ACM Press, New York, NY 10036, USA, 1989. ISBN 0-89791-299-3. LCCN QA75.5 .A1371 1989.

ACM:1990:CRS

[ACM90] Conference Record of the Seventeenth Annual ACM Symposium on Principles of Programming Languages: Papers Presented at the Symposium, San Francisco, CA, 17–19 January 1990. ACM Press, New York, NY 10036, USA, January 1990. ISBN 0-89791-343-4. LCCN QA76.7 .A15 1990.

Adams:1984:SC

[Ada84]  Russ Adams.  String com-

puter. Robotics Age, 6(11):22, 24–25, November 1984. CODEN ROAGD2. ISSN 0197-1905.

Adriaens:1995:SEG

[Adr95] G. Adriaens. Simplified English grammar and style correction in an MT framework: the LRE SECC project. ASLIB Proceedings, 47(3):73–82, March 1995. CODEN ASLPAO. ISSN 0001-253X (print), 1758-3748 (electronic).

Arbabi:1994:AAN

[AFCB94] M. Arbabi, S. M. Fischthal, V. C.

Cheng, and E. Bart. Algorithms for Arabic name transliteration. IBM Journal of Research and Development, 38(2): 183–193, March 1994. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

URL http://www.almaden.ibm.
com/journal/rd38-2.html.

Angell:1983:ASC

[AFW83] Richard C. Angell, George E. Freund, and Peter Willett. Automatic spelling correction using a trigram similarity measure. Information Processing and Management, 19(4):255–261, ???? 1983. CODEN IPMADK. ISSN 0306-4573 (print), 1873-5371 (electronic).

Al-Hussaini:2017:EIB

[AH17] Leena Al-Hussaini. Experience:

Insights into the benchmarking data of Hunspell and Aspell spell checkers. Journal of Data and Information Quality (JDIQ), 8(3– 4):13:1–13:??, July 2017. CO-DEN ???? ISSN 1936-1955.

Abramovich:1981:FIM

[AHZ81] Israel Abramovich, Shimon Hochbaum, and Reuven Zilber. First intelligent microprocessored dictionary that corrects spelling errors. In Baal-Schem et al. [BS^{+}81], page 461. ISBN ???? LCCN TK 5 M42 1981. IEEE Catalog number 81CH1659-2.

Akin:1993:SCC

[AKGH93] H. L. Akin, S. Kuru, T. Guengor, and I. Hamzaoglu. A spelling checker and corrector for Turkish. In Anonymous [Ano93e], pages 113–120. ISBN 975-518-024-9. LCCN ????

Alberga:1967:SSM

[Alb67] Cyril N. Alberga. String similarity and misspellings. Communications of the Association for Computing Machinery, 10 (5):302–313, May 1967. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Altman:1990:VSC

[Alt90] R. Altman. VP-to-the-max: a spell checker for Ventura — and more. PC Publishing, 5(9):35–38, 1990. ISSN 0896-8209.

Amsler:1986:CLR

[Ams86] Robert A. Amsler. Computational lexicology: a research program. In Addie Mattox, editor, AFIPS Conference Proceedings; vol. 55 1986 National Computer Conference, pages 397–403. AFIPS Press, Arlington, VA, USA, 1986. ISBN 0-88283-049-X. LCCN ????

Alkhatib:2020:DLA

[AMS20] Manar Alkhatib, Azza Abdel Monem, and Khaled Shaalan. Deep learning for Arabic error detection and correction. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(5):

71:1–71:13, August 2020. CO-DEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/
10.1145/3373266.

Angus:1990:JGS

[Ang90] Jeff Angus. Jake Geller’s Spell-Code. Computer Language Magazine, 7(6):105–??, June 1990. CO-DEN COMLEF. ISSN 0749-2839.

Anonymous:1984:PMA

[Ano84] Anonymous. Publication Manual of the American Psychological Association. American Psychological Association, Washington, DC, USA, third edition, 1984. ISBN 0-912704-57-8. 208 pp. LCCN BF76.7 .P83 1983.

Anon:1985:DDS

[Ano85a] Anon. Dynamic dictionary supplement for spell check tasks. IBM Technical Disclosure Bulletin, 28(6):2334–??, November 1985. CODEN IBMTAA. ISSN 0018-8689.

Anonymous:1985:A

[Ano85b] Anonymous. Advertisement. InfoWorld, 7(44):28, November 4 1985. CODEN INWODU. ISSN 0199-6649.

Anon:1986:CES

[Ano86a] Anon. Concurrent edit and spell check in a computer system having multiple processors. IBM Technical Disclosure Bulletin, 29 (1):396–398, June 1986. CODEN IBMTAA. ISSN 0018-8689.

Anon:1986:GUS

[Ano86b] Anon. Generation of ‘user specified’ DOS fully qualified file specification. IBM Technical Disclosure Bulletin, 28(9):4199–4201, February 1986. CODEN IBM-TAA. ISSN 0018-8689.

Anon:1986:MDT

[Ano86c] Anon. Multi-language dictionaries for text processing. IBM Technical Disclosure Bulletin, 28(10): 4310–4311, March 1986. CODEN IBMTAA. ISSN 0018-8689.

Anon:1986:USS

[Ano86d] Anon. User-specified spelling supplement support. IBM Technical Disclosure Bulletin, 28(8): 3324–3325, January 1986. CO-DEN IBMTAA. ISSN 0018-8689. Anonymous:1989:NOM

[Ano89] Anonymous, editor. National Online Meeting proceedings–1989, New York, May 9–11, 1989. Learned Information, Medford, NJ, USA, 1989. ISBN 0-938734-34-2. LCCN QA76.55 .N37 1989.

Anonymous:1991:SAS

[Ano91] Anonymous. Speller: an alternative spelling correction system. In Presperin [Pre91], pages 103–104. ISBN 0-932101-29-1. ISSN 0883-4741. LCCN ????

Anonymous:1992:AIE

[Ano92a] Anonymous, editor. Artificial intelligence, expert systems, natural language: 12th International conference — June 1992, Avignon, France, Artificial Intelligence Expert Systems Natural Language — International Conference. EC2, Nanterre, France, 1992. ISBN 2-906899-72-0, 2-906899-75-5. LCCN ???? FF1850. Four volumes.

Anonymous:1992:CIC

[Ano92b] Anonymous, editor. COLING-92: 15th International conference on computational linguistics — August 1992, Nantes, France, volume 1 of Coling — Conference. ICCL, ????, 1992. ISBN 952-90-2028-7.

Anonymous:1992:IPW

[Ano92c] Anonymous. An integrated predictive wordprocessing and spelling correction system. In Presperin [Pre92], pages 369–370. ISBN 0-932101-30-5. ISSN 0883-4741. LCCN ????

Anonymous:1992:RLD

[Ano92d] Anonymous. Reclaiming lost disk space; extending computer life; and spelling-checker problems. Byte Magazine, 17(8):295– ??, August 1992. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1993:CSC

[Ano93a] Anonymous, editor. Control systems and computer science: 9th International conference — 25– 28 May 1993, Bucharest, International Conference on Control Systems and Computer Science. Institutul Politehnic Bucuresti. Dept. of Control and Computers, Bucuresti, Romania, 1993. Two volumes.

Anonymous:1993:ECA

[Ano93b] Anonymous, editor. EACL ’93: 6th Conference — April 1993, Utrecht, The Netherlands, Conference — European Chapter of the Association for Computational Linguistics. Association for Computational Linguistics, ????, 1993. ISBN 90-5434-014-2. LCCN ????

Anonymous:1993:IAS

[Ano93c] Anonymous, editor. IEEE/SEMI Advanced Semiconductor Manufacturing Conference and Workshop. Theme: Factory of the Future. ASMC ’93 Proceedings: October 18–19, 1993, Boston, Massachusetts. Semicond. Equipment and Mater. Int, Mountain View, CA, USA, 1993. ISBN 0-7803-1367-4 (casebound), 0-7803-1366-6 (softbound), 0-7803-1368-2 (microfiche). LCCN TK 7836 I42 1993. IEEE catalog number 93CH3337-3.

Anonymous:1993:LSC

[Ano93d] Anonymous. Legal-specific spell checking. The Lawyer’s PC, 11 (6):6, December 1993. ISSN 0740-0942.

Anonymous:1993:TAI

[Ano93e] Anonymous, editor. Turkish artificial intelligence and artificial neural network: 2nd Symposium — June 1993, Istanbul, Turkey, Ikinci Turk Yapay Zeka Ve Yapay Sinir Aglari Sempozyumu. Univ Bilgisayar, Istanbul, Turkey, 1993. ISBN 975-518-024-9. LCCN ????

Anonymous:1994:FCA

[Ano94a] Anonymous, editor. Fourth Conference on Applied Natural Language Processing: Association for Computational Linguistics: proceedings of the conference, 13–15 October 1994, Stuttgart, Germany, Conference on Applied Natural Language Processing. Association for Computational Linguistics, Morristown, NJ, USA, 1994.

Anonymous:1994:IIC

[Ano94b] Anonymous, editor. ICSLP 94. 1994 International Conference on Spoken Language Processing. Acoustical Soc. Japan, Tokyo, Japan, 1994 4 vol. 1994. 4 vol.

Anonymous:1994:PPP

[Ano94c] Anonymous, editor. PRICAI-94. Proceedings of the 3rd Pacific Rim International Conference on Artificial Intelligence. Int. Acad. Publishers, Beijing, China, 1994. Two volumes.

Anonymous:1994:SI

[Ano94d] Anonymous. Spell it 3. CD-ROM World, 9(11):86, December 1994. CODEN CDWOEV. ISSN 1066-274X.

Anonymous:1995:SCR

[Ano95] Anonymous. Spell check reject.

Off our backs, 25(6):14, June 1995. ISSN 0030-0071.

Arceneaux:1983:PMS

[Arc83] C. D. Arceneaux. A poor man’s spelling checker. Creative Comput, 9(6):228–236, June 1983.

Arnold:1983:DD

[Arn83] T. A. Arnold. Dictionary disk. A+, 1(2):142–144, December 1983. CODEN AAAAE6. ISSN 0740-1590.

Artigas:1989:CMS

[Art89] Robert Artigas, Jr. CUG248 — micro spell. C Users Journal, 7 (2):104–??, February 1989. ISSN 0898-9788.

Asimov:1982:QS

[Asi82] I. Asimov. A question of spelling. Popular Computing, 1(9):106– 107, July 1982. ISSN 0279-4721.

ASPRS:1991:GLP

[ASP91] ASPRS, editor. GIS-LIS ’91 proceedings: 28 October–1 November, The Inforum, Atlanta, Georgia. American Society for Photogrammetry and Remote Sensing, American Congress on Surveying and Mapping, Bethesda, MD, 1991. ISBN 0-944426-75-1. LCCN G70.2 .G57 1991. Two volumes.

ATT:1986:AUS

[AT&T86] AT&T. AT&T UNIX System Readings and Applications, volume I. Prentice-Hall, Inc., Englewood Cliffs, NJ 07632, USA, 1986. ISBN 0-13-938532-0. xiv + 397 pp. LCCN QA76.76.O63 U553 1986.

Berghel:1988:TPE

[BA88] H. Berghel and C. Andreu. TAL-ISMAN: a prototype expert system for spelling correction. In ACM SIGSMALL ’88 [ACM88], pages 107–113. LCCN QA 76.6 A28 1988.

Bub:1985:WWA

[BCK85] D. Bub, A. Cancelliere, and A. Kertesz. Whole-word and analytic translation of spelling-to-sound in a non-semantic reader. In bub [PMC85], pages xxii + 544. ISBN 0-8367-7026-9. LCCN RC394.W6 S971 1985, WL 340.6 S961 1985. UK£29.95.

Bass:1995:SNN

[BD95] S. Bass and S. Dunn. Software for next to nothing. PC World, 13 (5):118–124, 126, May 1995. CO-DEN PCWDDV. ISSN 0737-8939 (print), 1944-9143 (electronic).

Boian:1994:AWP

[BDT94] E. Boian, A. Danilchenco, and L. Topal. Automation of wordforming process in the Romanian language. Studies in Informatics and Control, 3(1):43–52, March 1994. ISSN 1220-1766.

Bentley:1985:PPS

[Ben85] J. L. Bentley. Programming pearls: a spelling checker. Communications of the Association for Computing Machinery, 28 (5):456–462, May 1985. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bentley:1986:PP

[Ben86] Jon Louis Bentley. Programming Pearls. Addison-Wesley, Reading, MA, USA, 1986. ISBN 0-201-10331-1 (paperback). viii + 195 pp. LCCN QA76.6.B453 1986. Reprinted with corrections.

Berghel:1986:SVP

[Ber86a] H. Berghel. Spelling verification in Prolog. SIGPLAN Notices, 21 (1):19–27, January 1986. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Berghel:1986:ECW

[Ber86b] Hal L. Berghel. Extending the capabilities of word processing software through Horn clause lexical databases. AFIPS Conference Proceedings, 55:251–257, 1986. CODEN AFPGBT. ISBN 0-88283-049-X. ISSN 0095-6880.

Berghel:1987:LFC

[Ber87] H. L. Berghel. A logical framework for the correction of spelling errors in electronic documents. Information Processing and Management, 23(5):477–494, September 1987. CODEN IPMADK. ISSN 0306-4573 (print), 1873-5371 (electronic).

Besling:1994:HSM

[Bes94] S. Besling. Heuristical and statistical methods for grapheme-to-phoneme conversion. In Trost [Tro94], pages 23–31.

Beyls:1988:ASF

[Bey88] Pascal Beyls. An adaptation of spell to French. In EUUG Newsletter, volume 8(1), pages 11–14. European UNIX Users Group, Buntingford, Herts, UK, Spring 1988.

Beddoes:1984:PAL

[BG84] M. P. Beddoes and H. Garudadri. Possible Application of LPC: Spelling Correction. Canadian Medical & Biological Engineering Soc, Gloucester, Ont, Can, 1984. ISBN 0-919529-18-6. 116–117 pp. LCCN ????

Ben-Hamadou:1992:SCA

[BH92] A. Ben-Hamadou. Spelling correction of Arabic texts from a robust affix analysis of the affected strings. In Anonymous [Ano92a], pages 55–68. ISBN 2-906899-72-0, 2-906899-75-5. LCCN ???? FF1850. Four volumes.

Betz:1995:LMS

[BH95] M. Betz and H. Hild. Language models for a spelled letter recognizer. In IEEE [IEE95], pages 856–859 vol.1. ISBN 0-7803-2431-5. LCCN TK 7882 S65 I16 1995. Five volumes. IEEE Catalog number 95CH35732.

Bickel:1987:ACM

[Bic87] Michael Allen Bickel. Automatic correction to misspelled names: a fourth-generation language approach. Communications of the Association for Computing Machinery, 30(3):224–228, March 1987. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Belkin:1992:SPF

[BIP92] Nicholas Belkin, Peter Ingwersen, and Annelise Mark Pejtersen, editors. SIGIR ’92: proceedings of the Fifteenth Annual International ACM SIGIR Conference on Research and Development in Information Retrieval: Copenhagen, Denmark, June 21–24, 1992. ACM Press, New York, NY 10036, USA, 1992. CO-DEN FASRDV. ISBN 0-89791-523-2. ISSN 0163-5840. LCCN QA76.9.D3I552 1992.

Borchardt:1994:CAS

[BJ94] F. L. Borchardt and E. M. T. Johnson, editors. CALICO ’94 Annual Symposium. Proceedings of the Computer Assisted Language Instruction Consortium. 1994 Annual Symposium. ‘Human Factors’. Duke Univ, Durham, NC, USA, 1994.

Baird:1986:COP

[BKP86] Henry S. Baird, Simon Kahan, and Theo Pavlidis. Components of an omnifont page reader. In IEEE [IEE86], pages 344–348. CODEN PICREG. ISBN 0-8186-0742-4. LCCN Q 327 I615 1986. IEEE catalog number 86CH2342-4.

Blair:1960:PCS

[Bla60] Charles R. Blair. A program for correcting spelling errors. Information and Control, 3(??):60–67, 1960.

Barth:1985:SSS

[BN85] Wilhelm Barth and Heinrich Nirschl. Sichere Sinnentsprechende Silbentrennung für die Deutsche Sprache. Angewandte Informatik, Applied Informatics, 27(4):152– 159, April 1985. CODEN AW-IFA7.

Bomberger:1982:PPS

[Bom82] Alan Bomberger. A poor person’s spelling checker. Dr. Dobb’s Journal of Software Tools, 7(4): 42–53, April 1982. CODEN DDJOEB. ISSN 1044-789X.

Bos:1994:EDT

[Bos94] Edwin Bos. Error diagnosis in a tutoring system for the conjugation and spelling of Dutch verbs. Computers in Human Behavior, 10(1):33–49, Spring 1994. CO-DEN CHBEEQ. ISSN 0747-5632.

Bourne:1977:FIS

[Bou77] Charles P. Bourne. Frequency and impact of spelling errors in bibliographic data bases. Information Processing and Management, 13(1):1–12, ???? 1977. CO-DEN IPMADK. ISSN 0306-4573 (print), 1873-5371 (electronic).

Boyanov:1985:NOA

[Boy85] Kiril Boyanov, editor. Networks in Office Automation: Proceedings of the IFIP TC 6 International In-Depth Symposium on Networks in Office Automation, Sofia, Bulgaria, 25–30 September, 1984. North-Holland Publishing Co., Amsterdam, The Netherlands, 1985. ISBN 0-444-87715-0. LCCN HF5547.5 .I34 1985.

Breen:1993:JED

[Bre93] J. W. Breen. A Japanese electronic dictionary project (part 1:

The dictionary files). Technical Report 93-13, Monash University, Clayton, Victoria, Australia, November 30 1993. 25 pp. Available via anonymous

ftp in pub/techrports/RDT/93-
13.ps.Z from ftp.rdt.monash.
edu.au.

Balakrishnan:1994:CSE

[BRSS94] N. Balakrishnan, T. Radhakrishnan, D. Sampath, and S. Sundaram, editors. Computer Systems and Education. Proceedings of the International Conference on Computer Systems and Education in Honour of Prof. V. Rajaraman. Tata McGraw-Hill, New Delhi, India, 1994. ISBN 0-07-462044-4. LCCN ????

Baal-Schem:1981:OAK

[BS^{+}81] J. Baal-Schem et al., editors. Electrotechnology for development: proceedings of MELE-CON ’81, the first Mediterranean Electrotechnical Conference, Tel-Aviv, Israel, 24–28 May, 1981. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1981. ISBN ???? LCCN TK 5 M42 1981. IEEE Catalog number 81CH1659-2.

Badre:1982:DHC

[BS82] Albert Badre and Ben Shneiderman, editors. Directions in Human-Computer Interaction, Human/Computer Interaction. Ablex Publishing Corp., Norwood, NJ, USA, 1982. ISBN 0-89391-144-5. LCCN QA76.9.I58 .D57 1982.

Burton:1983:SIC

[BS83] Sarah K. Burton and Douglas D.

Short, editors. Sixth International Conference on Computers and the Humanities (North Carolina State University, Raleigh, NC, June 6–8, 1983). Computer Science Press, 11 Taft Court, Rockville, MD 20850, USA, 1983. ISBN 0-914894-96-X. LCCN AZ105 .I56 1983.

Bakar:2000:ERE

[BSY00] Zainab Abu Bakar, Tengku Mohd T. Sembok, and Mohammed Yusoff. An evaluation of retrieval effectiveness using spelling-correction and stringsimilarity matching methods on Malay texts. Journal of the American Society for Information Science, 51(8):691–706, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Buchan:1993:VT

[Buc93] Ronald L. Buchan. Variant terminology. In Strehlow and Wright [SW93], pages 95–105. CODEN ASTTA8. ISBN 0-8031-1493-1. ISSN 0066-0558. LCCN T11.S677 1993. ASTM publication code number (PCN) 04-011660-42.

Buyuk:2020:CDS

[Büy20]  Osman Büyük.  Context-

dependent sequence-to-sequence Turkish spelling correction. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(4): 56:1–56:16, July 2020. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL

https://dl.acm.org/doi/abs/
10.1145/3383200.

Carroll:1989:MPB

[CA89] J. Carroll and J. Abaitua. A morphological parser for Basque verbs’ inflection. In Campbell and Cuena [CC89], pages 77– 85. ISBN 0-470-21434-1 (Halsted Press: v. 1), 0-470-21435-X (Halsted Press: v. 2), 0-7458-0659-7 (v. 1), 0-7458-0660-0 (v. 2). LCCN Q335 .P42 1989. US$29.95 (v. 1), US$34.95 (v. 2). From Computing Reviews: “The authors briefly describe feature-value pairs based on LFG, a rewrite grammar, spelling rules, feature specifications and co-occurrences restrictions (FCRs), and rewrite rules.”.

Cannon:1981:OAK

[Can81] Theodore W. Cannon. Optically actuated keyboard system. In Proceedings of the Johns Hopkins 1st National Search for Applications of Personal Computing to Aid the Handicapped, pages 186–189. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1981. IEEE Serv Cent. Piscataway, NJ, USA.

Carpenter:1993:GNU

[Car93] Bruce Carpenter. Genealogy notes: Using soundex alternatives: Enumeration districts, 1880-1920. Prologue: the journal of the National Archives, 25 (1):90, Spring 1993. ISSN 0033-1031.

Cavnar:1994:NTF

[Cav94] W. B. Cavnar. N-gram-based text filtering for TREC-2. In Harman [Har94b], pages 171–179. NIST-SP 500-215.

Caudill:1987:IFI

[CB87] Maureen Caudill and Charles Butler, editors. IEEE First International Conference on Neural Networks, Sheraton Harbor Island East, San Diego, California, June 21–24, 1987. SOS Print, Piscataway, NJ, USA, 1987. ISBN (none). LCCN QP363.3 .I4 1987 v. 1–4.

Camuse:1984:MEC

[CBC84] Ruth A. Camuse, Gary G. Bitter, and Donna Craighead, editors. Microcomputers in Education Conference: Literacy Plus (Arizona State University, Tempe, AZ, March 13–15, 1984). Computer Science Press, 11 Taft Court, Rockville, MD 20850, USA, 1984. ISBN 0-88175-077-8. LCCN LB1028.5 .M53 1984.

Campbell:1989:PAI

[CC89] J. A. Campbell and J. Cuena, editors. Perspectives in Artificial Intelligence: Collection of Invited and Contributed Papers Presented at the Conference on Artificial Intelligence, held Sept. 7 and 11, 1987 in San Sebastian, Spain, within the framework of the Second World Basque Congress, Ellis Horwood series in artificial intelligence.

Halsted Press, New York, USA, 1989. ISBN 0-470-21434-1 (Halsted Press: v. 1), 0-470-21435-X (Halsted Press: v. 2), 0-7458-0659-7 (v. 1), 0-7458-0660-0 (v. 2). LCCN Q335 .P42 1989. US$29.95 (v. 1), US$34.95 (v. 2). Two volumes.

Ciura:2001:HSL

[CD01] Marcin G. Ciura and Sebastian Deorowicz. How to squeeze a lexicon. Software—Practice and Experience, 31(11):1077–1090, ???? 2001. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://www3.

                                        interscience.wiley.com/cgi-
                                        bin/abstract/85004330/START;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=85004330&PLACEBO=IE.pdf.

Cojocary:1993:RSC

[CEU93] S. Cojocary, M. Evstiunin, and V. Ufnarovski. Romanian spelling checker. In Anonymous [Ano93a], pages 123–136. Two volumes.

Cojocaru:1994:RS

[CEU94] S. Cojocaru, M. Evstiunin, and V. Ufnarovski. Romanian spelling-checker. Studies in Informatics and Control, 3(1):53– 58, March 1994. ISSN 1220-1766.

Cole:1991:SIN

[CFGJ91] R. A. Cole, M. Fanty, M. Gopalakrishnan, and R. D. T. Janssen. Speaker-independent name retrieval from spellings using a database of 50,000 names. In ICASSP 91: 1991 International Conference on Acoustics, Speech, and Signal Processing, May 14– 17, 1991, The Sheraton Centre Hotel and Towers, Toronto, Ontario, Canada. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1991.

Church:1991:PSS

[CG91] Kenneth W. Church and William A.

Gale. Probability scoring for spelling correction. Statistics and Computing, 1(2):93–103, December 1991. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http:/

/link.springer.com/article/
10.1007/BF01889984.

Chang:1994:PSA

[Cha94]  Chao-Huang Chang.  A pi-

lot study on automatic Chinese spelling error correction. Communications of COLIPS, 4(2): 143–149, December 1994. CO-DEN CIPSE9. ISSN 0218-7019.

Cherry:1981:CAW

[Che81] L. Cherry. Computer aids for writers. In Cherry [ACM81], pages 61–67. ISBN 0-89791-043-5. LCCN QA76.7 .S54 v.16:6. ACM SIGPLAN Notices, v. 16, no. 6, (June 1981).

Cheatham:1987:MDS

[Che87] Tom Cheatham. Message decryption and spelling checkers. Cryptologia, 11(2):115–118, April 1987. CODEN CRYPE6. ISSN 0161-1194 (print), 1558-1586 (electronic).

Chin:1984:ASG

[Chi84] David Chin. Analysis of scripts generated in writing between users and computer consultants. AFIPS Conference Proceedings, 53:637–642, 1984. CODEN AF-PGBT. ISBN 0-88283-043-0. ISSN 0095-6880.

Cox:1994:NCT

[CHJ94] Kevin Cox, Robert Hoffman, and Ernest Jordan. New computer tool for writers and readers. IEEE International Professional Communications Conference, pages 272–277, 1994. IEEE catalog number 94CH3442-1.

Christian:1987:UTP

[Chr87] Kaare Christian. The UNIX text processing system. John Wiley and Sons, Inc., New York, NY, USA, 1987. ISBN 0-471-85581-2. xii + 250 pp. LCCN QA76.76.O63 C47 1987. US$24.95. From Computing Reviews: “The short closing chapter sketches the use of a few simple tools for checking spelling, checking formatting commands, and previewing output on a terminal. Seven appendices provide useful checklists of commands, options, special characters, and macros.”.

Caramazza:1987:RGB

[CMVR87] Alfonso Caramazza, Gabriele Miceli, Giampiero Villa, and Cristina Romani. The role of the graphemic buffer in spelling: Evidence from a case of acquired dysgraphia. Cognition, 26(1):59– 85, June 1987. CODEN CGT-NAU. ISSN 0010-0277 (print), 1873-7838 (electronic).

Cook:2005:HCE

[Coo05] Robert P. Cook. Heuristic compression of an English word list. Software—Practice and Experience, 35(6):577–581, May 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Church:1995:CAN

[CR95] Kenneth W. Church and Lisa F.

Rau. Commercial applications of natural language processing. Communications of the Association for Computing Machinery, 38(11):71–79, November 1995. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cai:2016:DQA

[CRD16] Fei Cai, Ridho Reinanda, and Maarten De Rijke. Diversifying query auto-completion. ACM Transactions on Information Systems, 34(4):25:1–25:??, September 2016. CODEN ATISET. ISSN 1046-8188.

Cole:1992:TSD

[CRF92] R. Cole, K. Roginski, and M. Fanty. A telephone speech database of spelled and spoken names. In Proceedings of the International Conference on Spoken Language Processing. ????, ????, 1992.

Cribbs:1988:WLS

[Cri88] Margaret A. Cribbs. What to look for in a spelling checker — an essential tool for the micro. Online, 12(5):45–48, September 1988. CODEN ONLIDN. ISSN 0146-5422.

Comer:1982:HBS

[CS82] Douglas Comer and Vincent Y. Shen. Hash-bucket search — a fast technique for searching an English spelling dictionary. Software—Practice and Experience, 12(7):669–682, July 1982. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Calha:1992:SBP

[CT92] M. J. Calha and I. C. Teixeira. Speller board for personal computer. In Proceedings Euro ASIC ’92, Paris, France June 1– 5, 1992, pages 250–255 (of xii + 423). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1992. ISBN 0-8186-2845-6. LCCN TK 7874.6 E97 1992.

Cunningham:1994:MII

[Cun94] George P. Cunningham. Multiculturalism: Is it the spell checker, or is it just me? The Journal of the Midwest Modern Language Association, 27(1):70, Spring 1994. ISSN 0742-5562.

Cherkassky:1990:CAM

[CVB90] Vladimir Cherkassky, Nikolaos Vassilas, and Gregory L. Brodt. Conventional and associative memory-based spelling checkers. In IEEE [IEE90], pages 138–144. ISBN 0-8186-2084-6. LCCN Q334 .I565 1990. IEEE catalog number 90CH2915-7.

Cha:1990:KCM

[CW90] Sang K. Cha and Gio Wiederhold. Kaleidoscope: a cooperative menu-guided query interface. In Garcia-Molina and Jagadish [GMJ90], page 387. ISBN 0-89791-365-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.19 no.2 1990; QA1 .A87. URL http:/

/www.acm.org/pubs/articles/
proceedings/mod/93597/p387-
cha/p387-cha.pdf; http://www.
acm.org/pubs/citations/proceedings/
mod/93597/p387-cha/.

Comeau:2004:NWI

[CW04] Donald C. Comeau and W. John Wilbur. Non-word identification or spell checking without a dictionary. Journal of the American Society for Information Science and Technology: JASIST, 55(2): 169–177, January 15, 2004. CO-DEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Chen:2015:PFC

[CWC15] Kuan-Yu Chen, Hsin-Min Wang, and Hsin-Hsi Chen. A probabilistic framework for Chinese spelling check. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 14(4):15:1–15:??, October 2015. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). Special issue on Chinese spell checking.

Daly:1987:RWT

[Dal87] N. Daly. Recognition of words from their spellings: Integration of multiple knowledge sources. Master’s thesis, Massachusetts Institute of Technology, May 1987.

Damerau:1964:TCD

[Dam64] Fred J. Damerau. A technique for computer detection and correction of spelling errors. Communications of the Association for Computing Machinery, 7(3): 171–176, March 1964. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Davidson:1962:RMN

[Dav62] Leon Davidson. Retrieval of misspelled names in an airline passenger record system. Communications of the Association for Computing Machinery, 5(3): 169–171, March 1962. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

DAlbis:2012:PSC

[DBT^{+}12] Tiziano D’Albis, Rossella Blatt, Roberto Tedesco, Licia Sbattella, and Matteo Matteucci. A predictive speller controlled by a brain-computer interface based on motor imagery. ACM Transactions on Computer-Human Interaction, 19(3):20:1–20:??, October 2012. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Dobing:1985:SCB

[DC85] B. R. Dobing and J. E. Cooke. Spelling correction based on user error patterns. Technical report, University of Saskatchewan (??), Saskatoon, SK, Canada, 1985.

Du:1992:MFA

[DC92] M. W. Du and S. C. Chang.

A model and a fast algorithm for multiple errors spelling correction. Acta Informatica, 29(3): 281–302, June 1992. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Du:1994:ADV

[DC94] M.-W. Du and S. C. Chang. Approach to designing very fast approximate string matching algorithms. IEEE Transactions on Knowledge and Data Engineering, 6(4):620–633, August 1994. CODEN ITKEEH. ISSN 1041-4347.

Devlin:1994:CMc

[Dev94] Keith Devlin. Computers and mathematics. Notices of the American Mathematical Society, 41(7):772–??, September 1994. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Deffner:1990:RWC

[DGK^{+}90] R. Deffner, H. Geiger, R. Kahler, T. Krempl, and W. Brauer. Recognizing words with connectionistic architectures. In INNC ’90 [INN90], page 196. ISBN 0-7923-0831-X. LCCN QA76.87 .I584 1990.

Dembitz:2009:S

[DGR09] Šandor Dembitz, Gordan Gledec, and Mirko Randić. Spellchecker.

In Benjamin W. Wah, editor, Wiley Encyclopedia of Computer Science and Engineering, page ?? John Wiley and Sons, Inc., New York, NY, USA, 2009. ISBN 0-471-38393-7 (set), 0-470-05012-8 (v. 1), 0-470-05013-6 (v. 2), 0-470-05014-4 (v. 3), 0-470-05015-2 (v. 4), 0-470-05016-0 (v. 5). LCCN QA76.15 .W545 2009.

Ditlea:1987:HWD

[Dit87] S. Ditlea. How well do you spell? Personal Computing, 11 (3):91–103, March 1987. CODEN PLCMDL. ISSN 0192-5490.

Durham:1983:SCU

[DLS83] Ivor Durham, David A. Lamb, and James B. Saxe. Spelling correction in user interfaces. Communications of the Association for Computing Machinery, 26 (10):764–773, October 1983. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

deHaan:1994:SRS

[dO94] Ab de Haan and Tinus Oppenhuizen. SPELLER. A reflexive ITS to support the learning of second language spelling. Computers in Human Behavior, 10 (1):21–31, Spring 1994. CODEN CHBEEQ. ISSN 0747-5632.

Dolgopolov:1986:ASC

[Dol86] A. S. Dolgopolov. Automatic spelling correction. Cybernetics (English Translation of Kibernetika), 22(3):332–339, May– June 1986. CODEN CYBNAW. ISSN 0011-4235.

Dolgopolov:1988:SDCa

[Dol88a] A. S. Dolgopolov. Sequential decoding in a channel with spelling errors. Problems of Information Transmission, 24(1):19–24, July 1988. CODEN PRITA9. ISSN 0032-9460 (print), 1608-3253 (electronic).

Dolgopolov:1988:SDCb

[Dol88b] A. S. Dolgopolov. Sequential decoding in a channel with spelling mistakes. Problemy Peredachi Informatsii, 24(1):25–42, ???? 1988. CODEN PPDIA5. ISSN 0555-2923.

Desmarais:1994:OCL

[DR94] L. Desmarais and J. R. Roy.

ORTHO-DIDAC: courseware for learning French spelling. In Borchardt and Johnson [BJ94], pages 75–77.

Dembitz:2011:AOS

[DRG11] Šandor Dembitz, Mirko Randić, and Gordan Gledec. Advantages of online spellchecking: a Croatian example. Software— Practice and Experience, 41(11): 1203–1231, October 2011. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

DasGupta:1994:SCB

[DS94] S. Das Gupta and V. Sridhar.

Spelling correction based on hidden Markov models. In Balakrishnan et al. [BRSS94], pages 163–173. ISBN 0-07-462044-4. LCCN ????

Dunlavey:1981:LES

[Dun81]  Michael R. Dunlavey.  Let-

ter to the Editor: On spelling correction and beyond. Communications of the Association for Computing Machinery, 24 (9):608, September 1981. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Pet80a, Mil81, Nix81, Pet81].

Ebihara:1992:KSC

[Ebi92] Yoshihiko Ebihara. Knowledgebased spelling correction in Unix command names. Journal of Information Processing, 15(3):394– 399, 1992. CODEN JIPRDE. ISSN 0387-6101.

Emirkanian:1987:DCE

[EBR87] Louisette Emirkanian, Lorne H. Bouchard, and Sylvie Ratte. La détection et la correction des erreurs d’orthographe: un exemple d’ingénieire de la connaissance. [English: Detecting and correcting spelling errors, example of knowledge engineering]. In IEEE [IEE87b], pages 48–51. ISBN ???? LCCN R856.A2 C66 1987. IEEE Service Cent. Piscataway, NJ, USA.

Edmonds:1982:SCI

[Edm82] James D. Edmonds, Jr. Speedspelling: Could it work in the twenty-first century? Speculations in Science and Technology, 5(4):375–378, October 1982. CO-DEN SPSTDD. ISSN 0155-7785.

Edwards:1987:RU

[Edw87] Jon Edwards. RAM-resident utilities. Byte Magazine, 12(7), Summer 1987. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Ehsan:2013:GCS

[EF13] Nava Ehsan and Heshaam Faili. Grammatical and contextsensitive error correction using a statistical machine translation framework. Software— Practice and Experience, 43(2): 187–206, February 2013. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ehrenfeucht:1988:NDM

[EH88] A. Ehrenfeucht and D. Haussler.

A new distance metric on strings computable in linear time. Discrete Applied Mathematics, 20 (3):191–203, July 1988. CO-DEN DAMADU. ISSN 0166-218X (print), 1872-6771 (electronic).

Ehrenreich:1982:AAS

[EP82] S. L. Ehrenreich and T. Porcu.

Abbreviations for automated systems: teaching operators the rules. In Badre and Shneiderman [BS82], pages 111–135. ISBN 0-89391-144-5. LCCN QA76.9.I58 .D57 1982.

ElSadany:1989:AMS

[ESH89] T. A. El-Sadany and M. A.

Hashish. An Arabic morphological system. IBM Systems Journal, 28(4):600–612, 1989. CO-DEN IBMSA7. ISSN 0018-8670.

Fidanova:1997:LAS

[Fid97] Stefka Fidanova. Linear array for spelling correction. Concurrency: Practice and Experience, 9(10):967–973, October 1997. CODEN CPEXEI. ISSN 1040-3108. URL http://www3.

                                        interscience.wiley.com/cgi-
                                        bin/abstract?ID=13811;   http:
                                        //www3.interscience.wiley.
                                        com/cgi-bin/fulltext?ID=13811&
                                        PLACEBO=IE.pdf.

Fraser:1990:LTE

[FK90] C. W. Fraser and B. Krishnamurthy. Live text (editing). Software—Practice and Experience, 20(8):851–858, August 1990. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Floyd:1987:HHS

[Flo87] Edwin T. Floyd. Hashing for high-performance searching. Dr. Dobb’s Journal of Software Tools, 12(2):34–??, February 1987. CO-DEN DDJTEQ. ISSN 0888-3076.

Flynn:1988:HPI

[Fly88] Laurie Flynn. … Hypercard programs includes spelling checker, utility, icon builder. InfoWorld, 10(7):32, February 15 1988. CO-DEN INWODU. ISSN 0199-6649.

Fournier:1994:SUW

[Fou94] J.-P. Fournier. Searching for unknown words with docile agents. In Trost [Tro94], pages 112–120.

Furuta:1989:IES

[FQA89] R. Furuta, V. Quint, and J. Andre. Interactively editing structured documents. Electronic Publishing—Origination, Dissemination, and Design, 1(1): 19–44, April 1989. CODEN EPODEU. ISSN 0894-3982.

Friedman:1992:TSE

[FS92] Carol Friedman and Robert Sideli. Tolerating spelling errors during patient validation. Computers and Biomedical Research, 25(5), October 1992. CO-DEN CBMRB7. ISSN 0010-4809 (print), 1090-2368 (electronic).

Furugori:1990:ISC

[Fur90] T. Furugori. Improving spelling checkers for Japanese users of English. IEEE Transactions on Professional Communication, 33(3): 138–142, September 1990. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Frantz:1982:DCH

[FW82] G. A. Frantz and R. H. Wiggins. Design case history: speak and spell learns to talk. IEEE Spectrum, 19(2):45–49, February 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fallside:1985:CSP

[FW85] Frank Fallside and William A.

Woods, editors. Computer speech processing. Prentice-Hall International, Englewood Cliffs, NJ 07632, USA, 1985. ISBN 0-13-163841-6. LCCN TK7895.S65 F351 1985. US$39.95.

Frisch:1988:SAC

[FZ88] Rudolf Frisch and Antonio Zamora. Spelling assistance for compound words. IBM Journal of Research and Development, 32 (2):195–200, March 1988. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Galletta:2005:DSC

[GDEJ05] Dennis F. Galletta, Alexandra Durcikova, Andrea Everard, and Brian M. Jones. Does spellchecking software need a warning label? Communications of the Association for Computing Machinery, 48(7):82–86, July 2005. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Goldenberg:1987:ELL

[GF87] Ernest Paul Goldenberg and Wallace Feurzeig. Exploring language with LOGO. Explorations in LOGO. MIT Press, Cambridge, MA, USA, 1987. ISBN 0-262-57065-3. xv + 338 pp. LCCN QA76.73.L63 G654 1987. US$19.95. From Computing Reviews: “As the foreword tells us, this is an introductory textbook on linguistics and is devoted to the study of structures and patterns in meaning, spelling, sound, and style. An outstanding novel contribution is its systematic utilization of the programming language LOGO as a pedagogical tool, which allows an active learning style that encourages a playful exploration. Well-chosen examples study generative grammars that describe gossip, jokes, poems, new words, business documents, and various other chunks of language. Of course, the main focus is on the English language, but there are also lots of illustrations from other well-known languages such as French, Italian, German and Spanish, and even from less-used ones such as Basque, Estonian, Polish, and Romanian.”.

Giangardella:1967:SCV

[GHR67] John J. Giangardella, James F.

Hudson, and Richard S. Roper. Spelling correction by vector representation using a digital computer. IEEE Transactions on Engineering Writing and Speech, EWS-10(2):57–62, December 1967. CODEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Garrison:1984:USS

[GHU84] D. A. Garrison, C. S. Himelstein, and R. J. Urquhart. Unfold skipping for spelling correction function. IBM Technical Disclosure Bulletin, 26(10B):5642– 5643, March 1984. CODEN IBM-TAA. ISSN 0018-8689.

Garcia-Molina:1990:PAS

[GMJ90] Hector Garcia-Molina and H. V.

Jagadish, editors. Proceedings of the 1990 ACM SIGMOD International Conference on Management of Data, May 23–25, 1990, Atlantic City, NJ, volume 19(2) of SIGMOD Record (ACM Special Interest Group on Management of Data). ACM Press, New York, NY 10036, USA, June 1990. ISBN 0-89791-365-5. ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA 76.9 D3 S53 v.19 no.2 1990; QA1 .A87.

Gibbs:1975:ESA

[GMW75] A. T. Gibbs, J. P. McCarthy, and D. Whelpton. Electronic spelling aid for use in speech therapy. Medical and Biological Engineering, 13(5):742–745, September 1975. CODEN MBENAU.

Gonnet:1983:UDB

[Gon83] Gaston H. Gonnet. Unstructured Data Bases or Very Efficient Text Searching. ACM Press, New York, NY 10036, USA, 1983. ISBN 0-89791-097-4. 117–124 pp. LCCN ???? ACM (order n 475830). Baltimore, MD, USA.

Gordon:1985:IIF

[Gor85] P. Gordon. An introduction to issues in foreign-language computer design. In Boyanov [Boy85], pages 127–134. ISBN 0-444-87715-0. LCCN HF5547.5 .I34 1985.

Greanias:1978:ASV

[GR78] E. C. Greanias and W. S. Rosebaum. Automatic spelling verification: Towards a system solution of the office. Tsement, ?? (??):225–231, ???? 1978.

Grimes:1983:DDE

[Gri83] Gary J. Grimes. Digital data entry glove interface device, November 8 1983. (Keywords: human interface, finger spelling).

Goodenough-Trepagnier:1984:ESP

[GT84] Cheryl Goodenough-Trepagnier.

Ecrire: System pour la communication non-verbale française en orthographe standard. In Ecrire: System Pour la Communication Non-Verbale Française en Orthographe Standard, pages 343– 344. Rehabilitation Engineering Soc of North America, Bethesda, MD, USA, 1984.

Galli:1968:ESC

[GY68] Enrico J. Galli and Hisao M.

Yamada. Experimental studies in computer-assisted correction of unorthographic text. IEEE Transactions on Engineering Writing and Speech, EWS-11(2):75–84, August 1968. CO-DEN IEEWAY. ISSN 0018-9405 (print), 2331-3706 (electronic).

Hart:1991:HRC

[Har91] Horace Hart. Hart’s Rules; For Compositors and Readers at the University Press, Oxford. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 39th edition, 1991. ISBN 0-19-212983-X. 182 pp. LCCN Z254 .H37 1983. UK£8.

Haralambous:1994:TK

[Har94a] Yannis Haralambous. Typesetting Khmer. Electronic Publishing—Origination, Dissemination, and Design, 7(4): 197–215, December 1994. CO-DEN EPODEU. ISSN 0894-3982.

Harman:1994:STR

[Har94b] D. K. (Donna K.) Harman, editor. Second Text REtrieval Conference (TREC-2). National Institute for Standards and Technology, Gaithersburg, MD, USA, 1994. NIST-SP 500-215.

Hsieh:2015:CCS

[HBHC15] Yu-Ming Hsieh, Ming-Hong Bai, Shu-Ling Huang, and Keh-Jiann Chen. Correcting Chinese spelling errors with word lattice decoding. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 14(4):18:1–18:??, October 2015. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). Special issue on Chinese spell checking.

Heinselman:1972:CDC

[Hei72] Russell Craig Heinselman. Computerized detection and correction of spelling errors in FOR-TRAN programs. Thesis (m.s.), University of Minnesota, Minneapolis, MN, USA, 1972. 76 pp.

Heideman:1991:ADN

[Hei91] Michael Heideman, T. Automated detection of naming errors in a street network database. In ASPRS [ASP91], pages 724– 729 (vol. 2). ISBN 0-944426-75-1. LCCN G70.2 .G57 1991. Two volumes.

Hess:1990:RFC

[Hes90] Thomas Hess, editor. 1990 Rochester FORTH Conference:

Embedded Systems, June 12– 16th, 1990, University of Rochester. Institute for Applied FORTH Research, Rochester, NY, USA, June 1990. ISBN 0-914593-10-

2. LCCN QA76.73.F24 R59 1990. US$35.00.

Hanna:1966:PCC

[HHHR66] Paul Robert Hanna, J. S. Hanna, R. E. Hodges, and E. H. Rudorf. Phoneme-Grapheme Correspondences as Cues to Spelling Improvement. U.S. Department of Health, Education, and Welfare, Washington, DC, USA, 1966. 1716 pp. LCCN PE1143.H3.

Howard:2020:PSA

[HI20] James P. Howard and II. Phonetic spelling algorithm implementations for R. Journal of Statistical Software, 95(??): ??, ???? 2020. CODEN JS-SOBK. ISSN 1548-7660. URL

https://www.jstatsoft.org/
index.php/jss/article/view/
v095i08;   https://www.jstatsoft.
org/index.php/jss/article/
view/v095i08/v95i08.pdf.

Hickey:1983:KSS

[Hic83] Bob Hickey. Kaypro steps up its software. Microcomputing, 7(12): 70–73, December 1983. CODEN MIRCDC.

Hatakeyama:1995:IEJ

[HK95] T. Hatakeyama and H. Kakuda.

The implementation and evaluation of a Japanese editor utilizing pronunciation-information. Transactions of the Information Processing Society of Japan, 36 (1):119–128, January 1995. CO-DEN JSGRD5. ISSN 0387-5806.

Hodges:1987:ASI

[Hod87] Richard E. Hodges. American spelling instruction: Retrospect and prospect. Visible Language, 21(2):??, April 1987. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://

                                        s3-us-west-2.amazonaws.com/
                                        visiblelanguage/pdf/V21N2_
                                        1987_E.pdf.

Hoenig:1988:WTP

[HP88] Alan Hoenig and Mitch Pfeffer. Writers’ tools I: PC spelling and grammar checkers. TUGboat, 9 (1):46–47, April 1988. ISSN 0896-3207.

Holden:1992:GTE

[HR92] E. J. Holden and G. G. Roy. The graphical translation of English text into Signed English in the hand sign translator system. Computer Graphics Forum, 11 (3):C357–C366, ???? 1992. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hiller:1993:SAS

[HRLJ93] S. Hiller, E. Rooney, J. Laver, and M. Jack. SPELL: An automated system for computeraided pronunciation teaching. Speech Communication, 13(??): 463–474, December 1993. CO-DEN SCOMDH. ISSN 0167-6393.

Hendrix:1978:DNL

[HSSS78] Gary G. Hendrix, Earl D. Sacerdoti, Daniel Sagalowicz, and Jonathan Slocum. Developing a natural language interface to complex data. ACM Transactions on Database Systems, 3 (2):105–147, June 1978. CO-DEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic). URL http://www.acm.
org/pubs/articles/journals/
tods/1978-3-2/p105-hendrix/
p105-hendrix.pdf;   http://
www.acm.org/pubs/citations/
journals/tods/1978-3-2/p105-
hendrix/.

Huang:1987:LEE

[Hua87] Nai-Kuan Huang. A learning experiment on English spelling rules. In Caudill and Butler [CB87], pages II:351–II:358. ISBN (none). LCCN QP363.3 .I4 1987 v. 1–4. IEEE Service Cent (catalog number 87TH0191-7). Piscataway, NJ, USA.

Hughes:2004:CSS

[Hug04] Elliott Hughes. Checking spelling in source code. SIGPLAN Notices, 39(12):32–38, December 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hurford:1987:EAT

[Hur87]  Jeanne M. Hurford.  Elec-

tronic authoring tools. In IEEE [IEE87a], pages 45–46. ISBN ???? LCCN T10.5.I2 1987. IEEE Service Cent. Piscataway, NJ, USA.

Huston:1981:VVC

[Hus81] Thomas J. Huston. VCATS 80: a visual computer augmented teaching system for the 80’s. In Proceedings of the Johns Hopkins 1st National Search for Applications for Personal Computing to Aid the Handicapped. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1981. IEEE Serv Cent. Piscataway, NJ, USA.

IEEE:1986:EIC

[IEE86] IEEE, editor. Eighth International Conference on Pattern Recognition, Paris, France, October 27–31, 1986: proceedings. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1986. CODEN PICREG. ISBN 0-8186-0742-4. LCCN Q 327 I615 1986. IEEE catalog number 86CH2342-4.

IEEE:1987:IEC

[IEE87a] IEEE, editor. IEEE engineering communication: a byte into the future: conference record / IPCC, International Professional Communication Conference, 87, October 14 to 16, 1987, Sheraton Hotel, Winnipeg, Canada. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1987. ISBN ???? LCCN T10.5.I2 1987. IEEE catalog number 87CH2428-1.

IEEE:1987:MCC

[IEE87b] IEEE, editor. Montech ’87 conferences: Conference sur les technologies biomedicales, 10, 11, 12 novembre 1987, Montreal = Biomedical technologies conference, November 10, 11, 12, 1987. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1987. ISBN ???? LCCN R856.A2 C66 1987. IEEE Catalog number 87Ch2919-7.

IEEE:1988:IIC

[IEE88] IEEE, editor. IEEE International Conference on Neural Networks, Sheraton Harbor Island, San Diego, California, July 24– 27, 1988. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1988. LCCN QP 363.3 I58 1988. Two volumes. IEEE catalog number 88CH2632-8.

IEEE:1990:PII

[IEE90] IEEE, editor. Proceedings of the 2nd International IEEE Conference on Tools for Artificial Intelligence: Hyatt Hotel, Dulles International Airport, Herndon, VA, USA, November 6–9, 1990. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1990. ISBN 0-8186-2084-6. LCCN Q334 .I565 1990. IEEE catalog number 90CH2915-7.

IEEE:1992:EAP

[IEE92a] IEEE, editor. Euro ASIC ’92:

proceedings, CNIT, Paris, June 1–5, 1992. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1992. ISBN 0-8186-2845-6. LCCN TK7874.6 .E87 1992. IEEE catalog no. 92TH0442-4.

IEEE:1992:ISI

[IEE92b] IEEE, editor. IJCNN-91-Seattle, International Joint Conference on Neural Networks, July 8–12, 1991, Washington State Convention and Trade Center, Seattle, WA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1992. ISBN 0-7803-0164-1. LCCN QA76.87.I57 1991b. Two volumes. IEEE catalog number: 91CH3049-4.

IEEE:1993:IPI

[IEE93a] IEEE, editor. IJCNN ’93-Nagoya. Proceedings of 1993 International Joint Conference on Neural Networks. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-7803-1421-2. LCCN QA76.87 .I57 1993. Three volumes. IEEE Catalog number 93CH3353-0.

IEEE:1993:PSI

[IEE93b] IEEE, editor. Proceedings of the Second International Conference on Document Analysis and Recognition, October 20– 22, 1993, Tsukuba Science City, Japan. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-8186-4960-7 (paper), 0-8186-4961-5 (microfiche). LCCN TA1640.I57 1992. IEEE Catalog number 93TH0578-5.

IEEE:1993:SII

[IEE93c] IEEE, editor. Second IEEE International Conference on Fuzzy Systems: San Francisco, California, March 28–April 1, 1993. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-7803-0615-5. LCCN TJ212.2.I3249 1993. Two volumes. IEEE catalog number: 93CH3136-9.

IEEE:1995:ICA

[IEE95] IEEE, editor. 1995 International Conference on Acoustics, Speech, and Signal Processing. Conference Proceedings. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-7803-2431-5. LCCN TK 7882 S65 I16 1995. Five volumes. IEEE Catalog number 95CH35732.

Ito:1983:HFO

[IK83] Tetsuro Ito and Makoto Kizawa.

Hierarchical file organization and its application to similar-string matching. ACM Transactions on Database Systems, 8(3):410– 433, September 1983. CO-DEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (elec-

tronic). URL http://www.acm.
org/pubs/articles/journals/
tods/1983-8-3/p410-ito/p410-
ito.pdf;   http://www.acm.
org/pubs/citations/journals/
tods/1983-8-3/p410-ito/.

Iida:1994:MAK

[IN94] T. Iida and Y. Nakamura. A method to accept katakana variants. Transactions of the Information Processing Society of Japan, 35(11):2276–2282, November 1994. CODEN JS-GRD5. ISSN 0387-5806.

INNC:1990:IPI

[INN90] INNC 90 Paris: International Neural Network Conference, July 9–13, 1990, Palais des Congres, Paris, France, volume 1. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1990. ISBN 0-7923-0831-X. LCCN QA76.87 .I584 1990.

Iverson:1991:PVA

[Ive91] K. E. Iverson. A personal view of APL. IBM Systems Journal, 30(4):582–593, 1991. CODEN IBMSA7. ISSN 0018-8670.

Ito:1985:OHF

[IY85] T. Ito and C. T. Yu. Optimization of a hierarchical file organization for spelling correction. In ACM SIGIR ’85 [ACM85], pages 131–137. LCCN Z699.A1 I659 1985. US$23.00.

Imasaki:1993:FAN

[IYME93] N. Imasaki, T. Yamaguchi, D. Montgomery, and T. Endo. Fuzzy artificial network and its application to a command spelling corrector. In IEEE [IEE93c], pages 635–640. ISBN 0-7803-0615-5. LCCN TJ212.2.I3249 1993. Two volumes. IEEE catalog number: 93CH3136-9.

Jackel:1983:CSF

[Jac83] Manfred Jackel. Context sensitive formatting. SIGPLAN Notices, 18(2):65–68, February 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

James:1984:DD

[Jam84]  Geoffrey James.  Document

databases. Van Nostrand Reinhold Co., New York, NY, USA, 1984. ISBN 0-442-28185-4. xiii + 184 pp. LCCN QA76.9.D3 J35 1985. US$28.95.

Jain:2019:UES

[JJJT19] Amita Jain, Minni Jain, Goonjan Jain, and Devendra K. Tayal. “UTTAM”: an efficient spelling correction system for Hindi language based on supervised learning. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 18(1):8:1–8:??, January 2019. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/ft_
gateway.cfm?id=3264620.

Jouvet:1993:SSR

[JLMG93] D. Jouvet, A. Laine, J. Monne, and C. Gagnoulet. Speakerindependent spelling recognition over the telephone. Proceedings of the International Conference on Acoustics, Speech, and Signal Processing, 2:II–235–II–238, 1993. CODEN IPRODJ. ISBN 0-7803-0946-4. ISSN 0736-7791. IEEE catalog number 93CH3252-4.

Johnson:1981:MWS

[Joh81]  J. Johnson.  Making WP

smarter. Datamation, 29(11):58– 62, November 1981. CODEN DTMNAT. ISSN 0011-6963.

Jared:1991:DWI

[JS91] Debra Jared and Mark S. Seidenberg. Does word identification proceed from spelling to sound to meaning? Journal of Experimental Psychology: General, 120(4): 358–394, December 1991. CO-DEN JPGEDD. ISSN 0096-3445 (print), 1939-2222 (electronic).

Junqua:1995:NSD

[JVFM95] J.-C. Junqua, S. Valente, D. Fohr, and J.-F. Mari. An N-best strategy, dynamic grammars and selectively trained neural networks for real-time recognition of continuously spelled names over the telephone. In IEEE [IEE95], pages 852–855 vol.1. ISBN 0-7803-2431-5. LCCN TK 7882 S65 I16 1995. Five volumes. IEEE Catalog number 95CH35732.

Kim:1994:PMA

[KC94] Deok-Bong Kim and Key-Sun Choi. A predictive morphological analysis of Korean without backtracking. In Anonymous [Ano94c], pages 712–717 vol.2. Two volumes.

Kwon:2004:SKW

[KKC04] Hyuk-Chul Kwon, Mi-Young Kang, and Sung-Ja Choi. Stochastic Korean word-spacing with smoothing using Korean spelling checker. International Journal of Computer Processing of Oriental Languages (IJCPOL), 17(4):239– ??, December 2004. CODEN ???? ISSN 0219-4279.

Kernighan:1982:UDP

[KL82] B. W. Kernighan and M. E.

Lesk. UNIX document preparation. In Nievergelt et al. [NCNS82], pages 1–20. ISBN 0-444-86493-8. LCCN Z244 .D63 1982. US$46.50. From Computing Reviews: “… introduces ed,

nroff/troff, eqn, tbl, refer,

and pic, and touches on spell,

style, and diction.”.

Kline:19xx:CRL

[Klixx] E. A. Kline. Computer-aided review lessons in English grammar and spelling. In ????, pages 329–

332. ????, ????, 19xx.

Knuth:1992:LP

[Knu92] Donald E. Knuth. Literate Programming, volume 27 of CSLI Lecture Notes. Stanford University Center for the Study of Language and Information, Stanford, CA, USA, 1992. ISBN 0-937073-80-6 (paperback), 0-937073-81-4 (hardcover). xv + 368 pp. LCCN QA76.6.K644. URL http://

csli-www.stanford.edu/publications/
literate.html.

Kashyap:1984:SCU

[KO84] R. L. Kashyap and B. J. Oommen. Spelling correction using probabilistic methods. Pattern Recognition Letters, 2(3): 147–154, March 1984.

Kernighan:1981:STP

[KP81] Brian W. Kernighan and P. J. Plauger. Software Tools in Pascal. Addison-Wesley, Reading, MA, USA, 1981. ISBN 0-201-10342-7. ix + 366 pp. LCCN QA76.6 .K493. US$13.95.

Kahan:1987:RPC

[KPB87a] S. Kahan, T. Pavlidis, and H. S. Baird. On the recognition of printed characters of any font and size. IEEE Transactions on Pattern Analysis and Machine Intelligence, PAMI-9(2):274–288, March 1987. CODEN ITPIDJ. ISSN 0162-8828.

Khan:1987:RPC

[KPB87b] Simon Khan, Theo Pavlidis, and Henry S. Baird. On the recognition of printed characters of any font and size. IEEE Transactions on Pattern Analysis and Machine Intelligence, PAMI-9(2):274–288, March 1987. CODEN ITPIDJ. ISSN 0162-8828.

Krantz:1984:R

[Kra84] Donald G. Krantz. RESORT. Dr. Dobb’s Journal of Software Tools, 9(11):88–93, November 1984. CODEN DDJSDM. ISSN 1044-789X.

Kreiner:1992:RTM

[Kre92] David S. Kreiner. Reaction time measures of spelling: Testing a two-strategy model of skilled spelling. Journal of Experimental Psychology: Learning, Memory, and Cognition, 18(4): 765–776, July 1992. CODEN JEPCEA. ISSN 0278-7393.

Kukich:1988:BPT

[Kuk88] Karen Kukich. Back-propagation topologies for sequence generation. In IEEE [IEE88], pages 301– 308. LCCN QP 363.3 I58 1988. Two volumes. IEEE catalog number 88CH2632-8.

Kukich:1992:SCT

[Kuk92a] Karen Kukich. Spelling correction for the telecommunications network for the deaf. Communications of the Association for Computing Machinery, 35 (5):80–90, May 1992. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kukich:1992:TAC

[Kuk92b] Karen Kukich. Techniques for automatically correcting words in text. ACM Computing Surveys, 24(4):377–439, December 1992. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Krause:1981:LGN

[KW81] W. Krause and G. Willee. Lemmatizing German newspaper texts with the aid of an algorithm. Comput. Hum., 15(2):101– 113, August 1981. CODEN CO-HUAD. ISSN 0010-4817. From Computing Reviews: “Lemmatization is defined as ‘the (automatic) reduction of word-forms that are related by the morphology of their inflection to a normalized spelling.’ Compound words (typical for the German language) are treated as separate lemmata.”.

Knave:1987:WDU

[KW87] Bengt Knave and Per-Gunnar Wideback, editors. Work with Display Units 86: Selected Papers from the International Scientific Conference on Work With Display Units, Stockholm, Sweden, May 12–15, 1986. North-Holland Publishing Co., Amsterdam, The Netherlands, 1987. ISBN 0-444-70171-0. LCCN RC965.V53 I57 1986. US$85.

Liang:1993:STC

[LAS93] S. Liang, M. Ahmadi, and M. Shridhard. Segmentation of touching characters in printed document recognition. In IEEE [IEE93b], pages 569–572. ISBN 0-8186-4960-7 (paper), 0-8186-4961-5 (microfiche). LCCN TA1640.I57 1992. IEEE Catalog number 93TH0578-5.

Lavenier:1992:HPS

[Lav92]  Dominique Lavenier.  High

performance systolic chip for spelling correction. In IEEE [IEE92a], pages 381–384. ISBN 0-8186-2845-6. LCCN TK7874.6 .E87 1992. IEEE catalog no. 92TH0442-4.

Lavenier:1993:ISA

[Lav93] Dominique Lavenier. An integrated 2D systolic array for spelling correction. Integration, the VLSI journal, 15(1):97–111, July 1993. CODEN IVJODL. ISSN 0167-9260.

Lawson:1985:TTT

[Law85] Veronica Lawson, editor. Tools for the Trade: Translating and the Computer 5: Proceedings of a Conference Jointly Sponsored by Aslib, the Association for Information Management, the Aslib Technical Translation Group, the Translators’ Guild, with the cosponsorship of the Commission of the European Communities: 10– 11 November 1983. Aslib, London, UK, 1985. ISBN 0-85142-180-6. LCCN P308.T6 1985. UK£19.00.

Lawrence:1992:CMV

[Law92] Daniel Lawrence. CUG374 — MicroSpell v2.0. C Users Journal, 10(12):121–??, December 1992. ISSN 0898-9788.

Loosemore:1992:NNM

[LBW92] Richard P. W. Loosemore, Gordon D. A. Brown, and Frances L. Watson. A neural net model of normal and dyslexic spelling. In IEEE [IEE92b], pages 231– 236. ISBN 0-7803-0164-1. LCCN QA76.87.I57 1991b. Two volumes. IEEE catalog number: 91CH3049-4.

Lewis:2003:DVS

[LC03] J. R. Lewis and P. M. Commarford. Developing a voicespelling alphabet for PDAs. IBM Systems Journal, 42(4): 624–638, ???? 2003. CO-DEN IBMSA7. ISSN 0018-8670.

URL http://www.research.
ibm.com/journal/sj/424/lewis.
pdf.

Liu:2016:SBM

[LC16] Liangliang Liu and Cungen Cao. A seed-based method for generating Chinese confusion sets. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 16(1):5:1– 5:??, December 2016. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Luo:2024:CJN

[LC24] Queenie Luo and Yung-Sung Chuang. Cleansing jewel: a neural spelling correction model built on Google OCR-ed Tibetan manuscripts. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 23(5):73:1–73:??, May 2024. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). URL https://dl.

acm.org/doi/10.1145/3654811.

Liu:2015:HRA

[LCDM15] Xiaodong Liu, Fei Cheng, Kevin Duh, and Yuji Matsumoto. A hybrid ranking approach to Chinese spelling check. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 14(4):16:1–16:??, October 2015. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). Special issue on Chinese spell checking.

Litecky:1976:SEE

[LD76] Charles R. Litecky and Gordon B. Davis. Study of errors, error-proneness, and error diagnosis in Cobol. Communications of the Association for Computing Machinery, 19(1):33–37, January 1976. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Li:2020:PQA

[LHF^{+}20] Ying Li, Jizhou Huang, Miao Fan, Jinyi Lei, Haifeng Wang, and Enhong Chen. Personalized query auto-completion for largescale POI search at Baidu maps. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 19(5): 70:1–70:16, August 2020. CO-DEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

URL https://dl.acm.org/doi/
10.1145/3394137.

Lefevre:1992:MMF

[LHR^{+}92] J.-P. Lefevre, S. M. Hiller, E. Rooney, J. Laver, and M.- G. Di Benedetto. Macro and micro features for automated pronunciation improvement in the SPELL system. Speech, 11(1):31– 44, March 1992.

Lucchesi:1993:AFA

[LK93] Claudio L. Lucchesi and Tomasz Kowaltowski. Applications of finite automata representing large vocabularies. Software—Practice and Experience, 23(1):15– 30, January 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lee:2015:ISI

[LLWL15] Lung-Hao Lee, Gina-Anne Levow, Shih-Hung Wu, and Chao-Lin Liu. Introduction to the special issue on Chinese spell checking. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 14(4):14:1–14:??, October 2015. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). Special issue on Chinese spell checking.

Lucassen:1984:ITA

[LM84] J. M. Lucassen and R. L. Mercer. Information theoretic approach to the automatic determination of phonemic baseforms. Proceedings of the International Conference on Acoustics, Speech, and Signal Processing, 3:42. 5. 1– 42. 5. 4, 1984. CODEN IPRODJ. ISSN 0736-7791. IEEE Service Cent. Piscataway, NJ, USA.

Lenman:1987:NEA

[LM87] S. Lenman and H. Marmolin. Naming errors and automatic error correction in humancomputer interaction. In Knave and Wideback [KW87], pages 838–846. ISBN 0-444-70171-0. LCCN RC965.V53 I57 1986. US$85.

Laxon:1992:NNF

[LMPK92] Veronica Laxon, Jackie Masterson, Maggie Pool, and Corriene Keating. Nonword naming: Further exploration of the pseudohomophone effect in terms of orthographic neighborhood size, graphemic changes, spellingsound consistency, and reader accuracy. Journal of Experimental Psychology: Learning, Memory, and Cognition, 18(4):730–748, July 1992. CODEN JEPCEA. ISSN 0278-7393.

Lee:1999:RTS

[LNL99] Kin Hong Lee, Mau Kit Michael Ng, and Qin Lu. Research: Text segmentation for Chinese spell checking. Journal of the American Society for Information Science, 50(9):751–759, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Liang:1995:RTC

[LO95] Rung-Huei Liang and Ming Ouhyoung. A real-time continuous alphabetic sign language to speech conversion VR system. Computer Graphics Forum, 14 (3):C/67–C/76, September 1995. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lewis:1981:CEP

[LT81] Bob Lewis and Donovan Tagg, editors. Computers in education: proceedings of the IFIP TC-3 3rd World Conference on Computers in Education, Lausanne, Switzerland, July 27–31, 1981. North-Holland Publishing Co., Amsterdam, The Netherlands, 1981. CODEN CEDUDS. ISBN 0-444-86255-2. LCCN LB2846.4.W67 1981.

Lyon:1978:PST

[Lyo78] Gordon Lyon. Packed scatter tables. Communications of the Association for Computing Machinery, 21(10):857–865, October 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mace:1987:RNU

[Mac87] Scott Mace. Revelation now uses proximity-scan toolkit: AI technology flags errors. InfoWorld, February 8 1987. CODEN IN-WODU. ISSN 0199-6649.

Morris:1974:CDT

[MC74] Robert Morris and Lorinda L. Cherry. Computer detection of typographical errors. Technical Memorandum 1019 (TM 73-1271-4), AT&T Bell Laboratories, Murray Hill, NJ, USA, April 3, 1974. ?? pp.

Morris:1975:ACT

[MC75a] Robert Morris and Lorinda L. Cherry. Appendix A. Common technical English words, and Appendix B. Possible typo’s and spelling errors. IEEE Transactions on Professional Communication, PC-18(1):57–64, March 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic). See [MC75b].

Morris:1975:CDT

[MC75b] Robert Morris and Lorinda L.

Cherry. Computer detection of typographical errors. IEEE Transactions on Professional Communication, PC-18 (1):54–56, March 1975. CO-DEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic). See supplemental word list [MC75a].

McCullough:1992:CU

[McC92] William McCullough. CUG360 — uspell. C Users Journal, 10(3): 118–??, March 1992. ISSN 0898-9788.

McIlroy:1982:DSL

[McI82]  M. Douglas McIlroy.  De-

velopment of a spelling list. IEEE Transactions on Communications, COM-30(1 pt 1): 91–99, January 1982. CO-DEN IECMBT. ISSN 0096-

1965. URL http://www.cs.
dartmouth.edu/~doug/spell.
pdf.

McIlroy:1995:DSL

[McI95] M. Douglas McIlroy. Development of a spelling list. Report, AT&T Bell Laboratories, Murray Hill, NJ 07974, March 11, 1995. 13 pp. URL https://

www.cs.dartmouth.edu/~doug/
spell.pdf.

McMahon:1986:STP

[MCM86] L. E. McMahon, L. L. Cherry, and R. Morris. Statistical text processing, pages 227–244. Volume I of ATT [AT&T86], 1986.

ISBN 0-13-938532-0. LCCN QA76.76.O63 U553 1986.

McWilliams:1982:W

[McW82] P. McWilliams. Word. Popular Computing, 1(5):120–121, March 1982. ISSN 0279-4721.

Mays:1991:CBS

[MDM91] Eric Mays, Fred J. Damerau, and Robert L. Mercer. Context based spelling correction. Information Processing and Management, 27(5):517, 1991. CO-DEN IPMADK. ISSN 0306-4573 (print), 1873-5371 (electronic).

Means:1987:CYC

[Mea87] Linda G. Means. Cn yur computr raed ths? Research Publication — General Motors Research Laboratories, December 1987. CO-DEN GMRLAW. ISSN 0099-9326.

Mor:1982:HCM

[MF82a] M. Mor and A. S. Fraenkel. A hash code method for detecting and correcting spelling errors. Communications of the Association for Computing Machinery, 25(12):935–940 (or 935– 938??), December 1982. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Mor:1982:REF

[MF82b] Moshe Mor and Aviezri S. Fraenkel. Retrieval in an environment of faulty texts or faulty queries. In Scheuermann [Sch82], pages 405–425. CODEN PICDD7. ISBN 0-12-624080-9. LCCN QA76.9.D3I558 1982. US$34.00.

Miller:1981:LES

[Mil81] Lance A. Miller. Letter to the Editor: On spelling correction and beyond. Communications of the Association for Computing Machinery, 24(9):608–609, September 1981. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Pet80a, Dun81, Nix81, Pet81].

Miller:1990:WNM

[Mil90] A. Richard Miller. What’s in a name? an MMSFORTH implementation of the russell-soundex method. In Hess [Hes90], pages 101–103. ISBN 0-914593-10-2. LCCN QA76.73.F24 R59 1990. US$35.00.

Miller:1995:RAC

[Mil95] John W. Miller. Random access from compressed datasets with perfect value hashing. IEEE International Symposium on Information Theory, pages 454–??, 1995. CODEN PISTFZ. ISSN 0271-4655. IEEE catalog number 95CB35738.

Mitchell:1985:SIE

[Mit85] H. Mitchell. Some IBM equipment for the translator’s office. In Lawson [Law85], pages 15– 20. ISBN 0-85142-180-6. LCCN P308.T6 1985. UK£19.00.

Mitton:1987:SCS

[Mit87a] Roger Mitton. Spelling checkers, spelling correctors and the misspellings of poor spellers. Information Processing and Management, 23(5):495–505, 1987. CO-DEN IPMADK. ISSN 0306-4573 (print), 1873-5371 (electronic).

Mitton:1987:SCC

[Mit87b] Roger Mitton. Spelling checkers,spelling correctors and the misspellings of poor spellers. Information Processing and Management, 23(5):495–505, September 1987. CODEN IPMADK. ISSN 0306-4573 (print), 1873-5371 (electronic).

Mullin:1990:TTS

[MM90] James K. Mullin and Daniel J. Margoliash. A tale of three spelling checkers. Software— Practice and Experience, 20(6): 625–630, June 1990. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

McAuley:1992:CAC

[MM92] S. M. McAuley and T. F. McLaughlin. Comparison of Add-A-Word and Compu Spell programs with low-achieving students. The Journal of Educational Research, 85(6):362, July 1992. CODEN JEDRAP. ISSN 0022-0671.

Mikami:1993:SCE

[MM93] I. Mikami and H. Maeda. Spelling check of English science papers using a software for text formatter. Engineering and Technology, 10(3):115–121, 1993. CO-DEN KTGIAR. ISSN 0454-1405.

Morgan:1969:SCS

[Mor69] H. L. Morgan. Spelling correction and systems programming. Technical Report TR 69-31, Cornell University, Ithaca, NY, USA, February 1969. (email lmc@cs.cornell.edu).

Morgan:1970:SCS

[Mor70] H. L. Morgan. Spelling correction in systems programs. Communications of the Association for Computing Machinery, 13 (3):90–94, March 1970. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Morgan:1982:NCC

[Mor82] Howard Lee Morgan, editor.

1982 National Computer Conference: June 7–10, 1982, Houston, Texas. AFIPS Press, Arlington, VA, USA, 1982. ISBN 0-88283-035-X. LCCN TK7885.A1 J6 1982.

Monsell:1992:LST

[MPG^{+}92] Stephen Monsell, Karalyn E.

Patterson, Andrew Graham, Claire H. Hughes, and Robert Milroy. Lexical and sublexical translation of spelling to sound: Strategic anticipation of lexical status. Journal of Experimental Psychology: Learning, Memory, and Cognition, 18(3):452–467, May 1992. CODEN JEPCEA. ISSN 0278-7393.

Morris:1974:WSH

[MT74] Robert Morris and Ken Thompson. Webster’s second on the head of a pin. Technical Memorandum 1082 (TM 74-1271-13), AT&T Bell Laboratories, Murray Hill, NJ, USA, July 15, 1974. ?? pp.

Marzal:1993:CNE

[MV93] Andres Marzal and Enrique Vidal. Computation of normalized edit distance and applications. IEEE Transactions on Pattern Analysis and Machine Intelligence, 15(9):926–932, September 1993. CODEN ITPIDJ. ISSN 0162-8828.

Manber:1994:AAM

[MW94] Udi Manber and Sun Wu. Algorithm for approximate membership checking with application to password security. Information Processing Letters, 50(4): 191–197, May 1994. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nievergelt:1982:DPS

[NCNS82] J. Nievergelt, G. Coray, J.-D. Nicoud, and A. C. Shaw, editors. Document Preparation Systems: A Collection of Survey Articles. Elsevier North-Holland, Inc., New York, NY, USA, 1982. ISBN 0-444-86493-8. xiv + 274 pp. LCCN Z244 .D63 1982. US$46.50.

Nicolson:1992:SRD

[NF92] R. I. Nicolson and A. J. Fawcett. Spelling remediation for dyslexic children using the selfspell programs. Lecture Notes in Computer Science, 602:503–??, 1992.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Newman:1984:SAB

[NH84] P. W. Newman and Barry R.

Horowitz. Spelling aid for the blind. Conference Proceedings — IEEE SOUTHEAST-CON, pages 558–562, 1984. CO-DEN CPISDM. ISSN 0734-7502. IEEE Service Cent. Piscataway, NJ, USA.

Nix:1981:ESE

[Nix81] Robert P. Nix. Experience with a space efficient way to store a dictionary. Communications of the Association for Computing Machinery, 24(5):297–298, 1981. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Nardini:2010:TSC

[NSV^{+}10] Franco Maria Nardini, Fabrizio Silvestri, Hossein Vahabi, Pedram Vahabi, and Ophir Frieder. On tag spell checking. In String processing and information retrieval, volume 6393 of Lecture Notes in Computer Science, pages 37–42. Springer-Ver-lag, Inc., Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010.

Oflazer:1992:PAW

[Ofl92] K. Oflazer. Parsing agglutinative word structures and its application to spelling checkers in Turkish. In Anonymous [Ano92b], pages 39–45. ISBN 952-90-2028-7.

Oflazer:1994:SCA

[OG94] K. Oflazer and C. Guezey. Spelling correction in agglutinative languages. In Anonymous [Ano94a], pages 194–195.

Oommen:1995:PRS

[OL95] B. J. Oommen and R. K. S. Loke. Pattern recognition of strings containing traditional and generalized transposition errors. Proceedings of the IEEE International Conference on Systems, Man and Cybernetics, 2:1154– 1159, 1995. CODEN PICYE3. ISSN 0884-3627. IEEE catalog number 95CB35767.

OShea:1984:AAI

[O’S84] Tim O’Shea, editor. Advances in artificial intelligence: proceedings of the Sixth European Conference on Artificial Intelligence, ECAI-84, Pisa, Italy, September 5–7, 1984. North-Holland Publishing Co., Amsterdam, The Netherlands, 1984. ISBN 0-444-87611-1. LCCN Q334.E97 1984.

Pain:1981:CAS

[Pai81] Helen Pain. Computer aid for spelling error classification in remedial teaching. In Lewis and Tagg [LT81], pages 297–302. CO-DEN CEDUDS. ISBN 0-444-86255-2. LCCN LB2846.4.W67 1981.

Parrott:1983:SDI

[Par83] R. D. Parrott. Spelling dictionary incorporating word association.

IBM Technical Disclosure Bulletin, 25(11B):6246–6247, April 1983. CODEN IBMTAA. ISSN 0018-8689.

Peterson:1980:CPD

[Pet80a] James Lyle Peterson. Computer programs for detecting and correcting spelling errors. Communications of the Association for Computing Machinery, 23(12): 676–687, December 1980. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remarks [Dun81, Mil81, Nix81, Pet81].

Peterson:1980:CPS

[Pet80b] James Lyle Peterson. Computer Programs For Spelling Correction. Springer-Verlag, Inc., Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. ISBN 3-540-10259-0. vi + 213 pp. LCCN QA76.6 .P474.

Peterson:1981:LES

[Pet81] James L. Peterson. Letter to the Editor: On spelling correction and beyond. Communications of the Association for Computing Machinery, 24(9):609, September 1981. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Pet80a, Dun81, Mil81, Nix81].

Peterson:1982:UWS

[Pet82] J. L. Peterson. Use of webster’s seventh collegiate dictionary to construct a master hyphenation list. In Morgan [Mor82], pages 665–670. ISBN 0-88283-035-X. LCCN TK7885.A1 J6 1982.

Peterson:1986:NUT

[Pet86] James L. Peterson. A note on undetected typing errors. Communications of the Association for Computing Machinery, 29 (7):633–637, July 1986. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Pfaffenberger:1986:SPC

[Pfa86]  Bryan Pfaffenberger.  The

scholar’s personal computing handbook: a practical guide. Little, Brown and Co., Boston, MA, USA, 1986. ISBN 0-316-70401-6. xiii + 359 pp. LCCN QA76.5 .P399 1986. US$15.95. From Computing Reviews: “Spelling checkers to help weed out silly typographical errors, indexing programs to help prepare what may be the most important part of any major written academic work, and so forth, are very well described. Unfortunately, this book has a very skimpy index! It does not provide an alphabetized list naming each of the programs mentioned in the book. One has to hunt for various index terms under which specific programs may be named.”.

Prindle:1987:CST

[PM87] Lucelle Prindle and T. F. McLaughlin. A computer spelling testing program with clerical English students: An empirical evaluation. Journal of Computer-Based Instruction, 14(4):146– 149, Autumn 1987. CODEN JCOID8. ISSN 0098-597X.

Parvez:2013:OAH

[PM13] Mohammad Tanvir Parvez and Sabri A. Mahmoud. Offline Arabic handwritten text recognition: a survey. ACM Computing Surveys, 45(2):23:1–23:??, February 2013. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Punchimudiyanse:2017:AFW

[PM17] Malinda Punchimudiyanse and Ravinda Gayan Narendra Meegama. Animation of fingerspelled words and number signs of the Sinhala sign language. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 16(4):24:1–24:??, September 2017. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Patterson:1985:SDN

[PMC85] K. E. Patterson, J. C. Marshall, and M. Coltheart. Surface Dyslexia: Neuropsychological and Cognitive Studies of Phonological Reading. Erlbaum, New York, NY, USA, and London, UK, 1985. ISBN 0-8367-7026-9. xxii + 544 pp. LCCN RC394.W6 S971 1985, WL 340.6 S961 1985. UK£29.95.

Pollock:1982:SED

[Pol82] J. J. Pollock. Spelling error detection and correction by computer: Some notes and a bibliography. Journal of Documentation, 38(4): 282–291, December 1982. CO-DEN JDOCAS. ISSN 0022-0418 (print), 1758-7379 (electronic).

Pournelle:1984:R

[Pou84] J. Pournelle. On the road. Byte Magazine, 9(10):363–382, September 1984. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Presperin:1991:TNA

[Pre91] J. J. Presperin, editor. Technology for the nineties: 14th Annual conference — June 1991, Kansas City, MO, volume 11 of Proceedings of the RESNA Annual Conference. RESNA Press, ????, 1991. ISBN 0-932101-29-1. ISSN 0883-4741. LCCN ????

Presperin:1992:TCI

[Pre92] J. J. Presperin, editor. Technology for consumers: International conference — June 1992, Toronto, Canada, volume 12 of Proceedings of the RESNA Annual Conference. Resna Press, Washington, DC, USA, 1992. ISBN 0-932101-30-5. ISSN 0883-4741. LCCN ????

Prosise:1995:TNM

[Pro95] Jeff Prosise. Tutor — no matter how you spell it, soundex finds it. PC Magazine, 14(10):261, 1995. CODEN PCMGEP. ISSN 0888-8507.

Pisarn:2007:HBM

[PT07] C. Pisarn and T. Theeramunkong. An HMM-based method for Thai spelling speech recognition. Computers and Mathematics with Applications, 54(1):76–95, July 2007. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0898122107001812.

Pirkola:2007:FBI

[PTKJ07] Ari Pirkola, Jarmo Toivonen, Heikki Keskustalo, and Kalervo Järvelin. Frequency-based identification of correct translation equivalents (FITE) obtained through transformation rules. ACM Transactions on Information Systems, 26(1):2:1– 2:??, November 2007. CODEN ATISET. ISSN 1046-8188.

Pollock:1982:ASE

[PZ82] Joseph J. Pollock and Antonio Zamora. Automatic spelling error detection and correction in textual databases. Proceedings of the ASIS annual meeting, 19:236– 238, 1982. CODEN PAISDQ. ISBN 0-86729-038-2. ISSN 0044-7870.

Pollock:1983:CCS

[PZ83] J. J. Pollock and A. Zamora.

Collection and characterization of spelling errors in scientific and scholarly text. Journal of the American Society for Information Science, 34(1):51–58, January 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Pollock:1984:ASC

[PZ84a] J. J. Pollock and A. Zamora. Automatic spelling correction in scientific and scholarly text. Communications of the Association for Computing Machinery, 27 (4):358–368, April 1984. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Pollock:1984:SDD

[PZ84b] Joseph J. Pollock and Antonio Zamora. System design for detection and correction of spelling errors in scientific and scholarly text. Journal of the American Society for Information Science, 35(2):104–109, March 1984. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Radue:1983:DIS

[Rad83] Jon Radue. On the design of an interactive spelling dictionary for personal computers. In ACM PSC ’83 [ACM83], pages 197– 199. ISBN 0-89791-123-7. LCCN QA 76.5 S53 v.6 no.2. ACM (Order n 609830). Baltimore, MD, USA.

Ramshaw:1994:CRS

[Ram94] Lance A. Ramshaw. Correcting real-word spelling errors using a model of the problem-solving context. Computational Intelligence, 10(2):185–211, May 1994. CODEN COMIE6. ISSN 0824-7935 (print), 1467-8640 (electronic)d.

Rzevski:1994:AAI

[RAR94] G. Rzevski, R. A. Adey, and D. W. Russell, editors. Applications of artificial intelligence in engineering IX: proceedings of the ninth international conference, held in Pennsylvania, USA, 19th–21st July, 1994. Comput. Mech. Publications, Southampton, UK, 1994. ISBN 1-85312-284-X. LCCN TA345.I5674 1994.

Robbins:2005:CSS

[RB05] Arnold Robbins and Nelson H. F. Beebe. Classic Shell Scripting. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00595-4. xxii + 534 pp. LCCN QA76.76.O63 R633 2005.

URL http://www.oreilly.com/
catalog/shellsrptg/.  Also

available in Chinese [?], French [?], German [?], Japanese [?], and Polish [?] translations.

Richardson:1998:TCI

[Ric98] Marjorie Richardson. Take command: ispell: Spelling checker. Linux Journal, 46:??, February 1998. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Rondthaler:1978:SRO

[RL78] Edward Rondthaler and Edward J. Lias. Soundspel: a revised orthography of the English language. IEEE Transactions on Professional Communication, PC-21(1):25–29, March 1978. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Robinson:1983:EWM

[Rob83] D. Robinson. Electric Webster: much more than a spelling checker. Creative Comput, 9(11): 108–112, November 1983.

Robertson:1986:ATW

[Rob86] Barbara Robertson. The AI typist: Writing aid is fast and easy, but bug plagued. InfoWorld, 8 (2):35, January 13 1986. CODEN INWODU. ISSN 0199-6649.

Rosenthal:1982:SCC

[Ros82] Eric S. Rosenthal. Spelling checkers, compound words, and variant spellings. In USENIX-Summer’92 [USE82], pages 315–322.

Robinson:1981:ASC

[RS81] P. Robinson and D. Singer. Another spelling correction program. Communications of the Association for Computing Machinery, 24:296–297, 1981. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ringlstetter:2007:ATC

[RSM07] Christoph Ringlstetter, Klaus U. Schulz, and Stoyan Mihov. Adaptive text correction with Webcrawled domain-dependent dictionaries. ACM Transactions on Speech and Language Processing (TSLP), 4(4):9:1–9:??, October 2007. CODEN ???? ISSN 1550-4875.

Rickman:1973:SIO

[RW73] J. Rickman and W. E. Walden. Structures for an interactive online thesaurus. International Journal of Computer and Information Sciences, 2(2):115–127, June 1973. CODEN IJCIAH. ISSN 0091-7036.

Robertson:1992:SHW

[RW92] Alexander M. Robertson and Peter Willett. Searching for historical word-forms in a database of 17th-century English text using spelling-correction methods. In Belkin et al. [BIP92], pages 256– 265. CODEN FASRDV. ISBN 0-89791-523-2. ISSN 0163-5840. LCCN QA76.9.D3I552 1992.

Rytting:2011:SCD

[RZR^{+}11] C. Anton Rytting, David M. Zajic, Paul Rodrigues, Sarah C. Wayland, Christian Hettick, Tim Buckwalter, and Charles C. Blake. Spelling correction for dialectal Arabic dictionary lookup. ACM Transactions on Asian Language Information Processing, 10(1):3:1–3:??, March 2011. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Sengupta:1996:MPI

[SC96] P. Sengupta and B. B. Chaudhuri. Morphological processing of Indian languages for lexical interaction with application to spelling error correction. Sadhana — Academy Proceedings in Engineering Sciences, 21(pt 3): 363–380, June 1996. CODEN SAPSER. ISSN 0256-2499.

Schek:1975:TFK

[Sch75] H.-J. Schek. Tolerating fuzzyness in keywords by similarity searches. Technical report, University of Heidelberg (??), Heidelberg, Germany, 1975.

Scheuermann:1982:PSI

[Sch82] Peter Scheuermann, editor. Proceedings of the Second International Conference on Databases: Improving Usability and Responsiveness, June 22–24, 1982, the Hebrew University of Jerusalem, Israel. Academic Press, New York, NY, USA, 1982. CODEN PICDD7. ISBN 0-12-624080-9. LCCN QA76.9.D3I558 1982. US$34.00.

Schulz:1994:FSS

[Sch94] U. Schulz. Fault-tolerant structure of search processes in OPACs. ABI-Technik, 14(4):299– 310, 1994. CODEN ABITD7. ISSN 0720-6763.

Stancikova:1994:EKO

[SD94] Pavla Stancikova and Ingetraut Dahlberg, editors. Environmental Knowledge Organization and Information Management. Proceedings of the First European ISKO Conference, 14–16 Sept. 1994, Bratislava, Slovakia. IN-DEKS Verlag, Frankfurt/Main, Germany, 1994. ISBN 3-88672-600-2. ISSN 0946-9389. LCCN GE30.E97 1994.

Sennhauser:1993:IRA

[Sen93] René Sennhauser. Improving the recognition accuracy of text recognition systems using typographical constraints. Electronic Publishing—Origination, Dissemination, and Design, 6(3): 273–282, September 1993. CO-DEN EPODEU. ISSN 0894-3982.

Soo:2015:BCS

[SF15] Jason Soo and Ophir Frieder.

Brief communications: On searching misspelled collections. Journal of the Association for Information Science and Technology, 66(6):1294–1298, June 2015. CODEN ???? ISSN 2330-1643 (print), 2330-1643 (electronic).

Suraj:2011:RPC

[SGM11] M. G. Suraj, D. S. Guru, and S. Manjunath. Recognition of postal codes from fingerspelling video sequence. International Journal of Image and Graphics (IJIG), 11(1):21–41, January 2011. CODEN ???? ISSN 0219-4678.

Srihari:1983:IDK

[SHC83] S. N. Srihari, J. J. Hull, and R. Choudhari. Integrating diverse knowledge sources in text recognition. ACM Transactions on Office Information Systems, 1(1): 68–87, January 1983. CODEN ATOSDO. ISSN 0734-2047.

Sidorov:1979:AWS

[Sid79] A. A. Sidorov. Analysis of word similarity in spelling correction systems. Programming and Computer Software; translation of Programmirovaniye (Moscow, USSR) Plenum, 5(4):274–277, July–August 1979. CODEN PC-SODA. ISSN 0361-7688 (print), 1608-3261 (electronic). English translation of Russian article in Programmirovanie 1979, No.4, 65-68 (1979).

Sillery:1985:SIR

[Sil85] B. Sillery. Spelling it right: Webster’s new word spelling checker. Personal Computing, 9 (11):162, November 1985. CO-DEN PLCMDL. ISSN 0192-5490.

Smith:1983:UWW

[SK83] C. R. Smith and K. E. Kiefer. Using the Writer’s Workbench programs at Colorado State University. In Burton and Short [BS83], pages 672–684. ISBN 0-914894-96-X. LCCN AZ105 .I56 1983. From Computing Reviews: “The Writer’s Workbench programs described include: ORGANIZATION; DEVELOP-MENT; FINDBE; DICTION; SUGGEST; SPELL; PUNCTU-ATION; GRAMMAR; STYLE; PROSE; ABSTRACT; VAGUE-NESS; and CHECK. The latter two programs were developed at CSU for inclusion in the package. In combination, the thirteen programs produce some five to six pages of analysis to aid students in revision before their papers are handed in for marking.”.

Skousen:1982:ESP

[Sko82] Royal Skousen. English spelling and phonemic representation. Visible Language, 16(1):28– 38, January 1982. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

URL https://s3-us-west-2.
amazonaws.com/visiblelanguage/
pdf/V16N1_1982_E.pdf.

Seni:1996:GED

[SKS96] Giovanni Seni, V. Kripasundar, and Rohini K. Srihari. Generalizing edit distance to incorporate domain information: handwritten text recognition as a case study. Pattern Recognition, 29 (3):405–414, March 1996. CO-DEN PTNRA8. ISSN 0031-3203 (print), 1873-5142 (electronic).

Slade:1987:PLD

[Sla87] Stephen Slade. The T programming language: a dialect of LISP. Prentice-Hall, Inc., Englewood Cliffs, NJ 07632, USA, 1987. ISBN 0-13-881905-X. xiv + 425 pp. LCCN QA76.73.T18 S58 1987. US$21.33. From Computing Reviews: “The wellthought-out problems on interesting topics such as spelling correction, cryptography, lazy array space allocation, workspace editor, database management system, etc. are particularly pleasing.”.

Sippl:1980:ECD

[SM80] Charles J. Sippl and JoAnn Coffman Mayer. The Essential Computer Dictionary and Speller for Secretaries, Managers, and Office Personnel. Prentice-Hall, Inc., Englewood Cliffs, NJ 07632, USA, 1980. ISBN 0-13-284364-1 (hardcover), 0-13-284356-0 (paperback). v + 258 pp. LCCN QA76.15 .S514. US$13.95 (hardcover), US$6.95 (paperback).

Smith:1983:SDL

[Smi83] M. W. A. Smith. Stylometry: the detection of literary authorship. The Computer Bulletin, 2(35):8– 9, 11, March 1983.

Sekar:1990:SDS

[SPR90] R. C. Sekar, Shaunak Pawagi, and I. V. Ramakrishnan. Small domains spell fast strictness analysis. In ACM POPL ’90 [ACM90], pages 169–183. ISBN 0-89791-343-4. LCCN QA76.7 .A15 1990.

Sinha:1984:PCS

[SS84] R. M. K. Sinha and K. S. Sigh. A program for correction of single spelling errors in Hindi words. Journal of the Institution of Electronics and Telecommunication Engineers, 30(6):249– 251, November 1984. CODEN JI-ETAU. ISSN 0377-2063.

Sharma:2014:WPS

[SS14] Manoj Kumar Sharma and Debasis Samanta. Word prediction system for text entry in Hindi. ACM Transactions on Asian Language Information Processing, 13(2):8:1–8:??, June 2014. CODEN ???? ISSN 1530-0226 (print), 1558-3430 (electronic).

Stultz:1982:WPH

[Stu82] R. A. Stultz. The word processing handbook. Prentice-Hall, Inc., Englewood Cliffs, NJ 07632, USA, 1982. ISBN 0-13-963454-1, 0-13-963447-9 (paperback). 237 pp. LCCN HF5548.115 .S78. US$17.95.

Strehlow:1993:STB

[SW93] Richard A. Strehlow and Sue Ellen Wright, editors. Standardizing terminology for better communication: practice, applied theory, and results, volume 1166. ASTM, Philadelphia, PA, USA, 1993. CODEN ASTTA8. ISBN 0-8031-1493-1. ISSN 0066-0558. LCCN T11.S677 1993. ASTM publication code number (PCN) 04-011660-42.

Taylor:1986:WAM

[Tay86] Dave Taylor. Wordz that almost match. Computer Language Magazine, 3(11):47–59, November 1986. CODEN COMLEF. ISSN 0749-2839.

Teitelman:1984:TTC

[Tei84] W. Teitelman. A tour through cedar. IEEE Software, 1 (2):44–73, April 1984. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). From Computing Reviews: “Cedar reduces many good ideas to practice, including such philosophical dicta as The Law of Least Astonishment (user experience should correctly predict system behavior), Do What I Mean (interaction mistakes, e.g., spelling errors, should be automatically corrected in context), and Unlimited Undo (every action should be reversible).”.

Terpenning:1981:CTH

[Ter81] Joseph L. Terpenning. ‘computertutor’ for the handicapped. In Proceedings of the Johns Hopkins 1st National Search for Applications of Personal Computing to Aid the Handicapped, pages 74–75. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1981. IEEE Serv Cent. Piscataway, NJ, USA.

Thomas:1983:DDE

[TG83] R. Thomas and B. Gustafson. The design, development and evaluation of a low-cost computermanaged spelling system. AEDS J, 16(3):168–176, Spring 1983.

Thimbleby:1982:BUE

[Thi82] H. Thimbleby. Basic user engineering principles for display editors. In Williams [Wil82], pages 537–541. ISBN 0-444-86464-4. LCCN TK5105.5 .I57 1982.

Takahashi:1990:SCM

[TIAY90] H. Takahashi, N. Itoh, T. Amano, and A. Yamashita. A spelling correction method and its application to an OCR system. Pattern Recognition, 23(3-4):363– 377, 1990. CODEN PTNRA8. ISSN 0031-3203 (print), 1873-5142 (electronic).

Tanaka:1986:HSS

[TKS86] Eiichi Tanaka, Takahiro Kohashiguchi, and Kunihiko Shimamura. High speed string correction for OCR. In IEEE [IEE86], pages 340–343. CO-DEN PICREG. ISBN 0-8186-0742-4. LCCN Q 327 I615 1986. IEEE Service Cent. Piscataway, NJ, USA.

Terrell:1982:EEL

[TL82] C. D. Terrell and O. Linyard.

Evaluation of electronic learning aids: Texas Instruments’ “Speak Spell”. International Journal of Man-Machine Studies, 17(1): 59–67, July 1982. CODEN IJMMBC. ISSN 0020-7373.

Trost:1994:KVN

[Tro94] H. Trost, editor. KONVENS ’94.

Verarbeitung Naturlicher Sprache (KONVENS ’94. Natural Speech Processing). Univ. Wien, Vienna, Austria, 1994.

Tsunoda:1993:CRA

[TST93] T. Tsunoda, T. Shiraishi, and H. Tanaka. Character recognition by associative completion on words. In IEEE [IEE93a], pages 1135–1138 vol.2. ISBN 0-7803-1421-2. LCCN QA76.87 .I57 1993. Three volumes. IEEE Catalog number 93CH3353-0.

Tanaka:1986:HSE

[TTK86] E. Tanaka, T. Toyama, and S. Kawai. High speed error correction of phoneme sequences. Pattern Recognition, 19(5):407– 412, 1986. CODEN PTNRA8. ISSN 0031-3203 (print), 1873-5142 (electronic).

Turba:1981:CST

[Tur81] T. N. Turba. Checking for spelling and typographical errors in computer-based text. In ACM SIGOA ’81 [ACM81], pages 101– 113 (or 51–60??). ISBN 0-89791-043-5. LCCN QA76.7 .S54 v.16:6. ACM SIGPLAN Notices, v. 16, no. 6, (June 1981).

Turba:1982:LSL

[Tur82] T. N. Turba. Length-segmented lists. Communications of the Association for Computing Machinery, 25(8):522–526, August 1982. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Trenkle:1994:DSC

[TV94] John M. Trenkle and Robert C. Vogt, III. Disambiguation and spelling correction for a neural network-based character recognition system. In Vincent and Pavlidis [VP94], pages 322–333. ISBN 0-8194-1476-X. ISSN 0361-0748. LCCN TA1640.D63 1994.

Teague:1984:UCI

[TWT84] G. V. Teague, R. M. Wilson, and M. G. Teague. Use of computerassisted instruction to improve spelling proficiency of low achieving first graders. AEDS J, 17(4): 30–35, Summer 1984.

Ullmann:1977:BGT

[Ull77] J. R. Ullmann. A binary ngram technique for automatic correction of substitution, deletion, insertion and reversal errors in words. The Computer Journal, 20(2):141–147, May 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.
co.uk/computer_journal/hdb/
Volume_20/Issue_02/tiff/141.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
20/Issue_02/tiff/142.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/143.tif;   http://
www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_02/tiff/
144.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_20/Issue_02/tiff/145.
tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_
20/Issue_02/tiff/146.tif;
http://www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_
02/tiff/147.tif.

Urquhart:1984:FSC

[Urq84] R. J. Urquhart. Filter for spelling correction function. IBM Technical Disclosure Bulletin, 26(10B): 5637–5638, March 1984. CODEN IBMTAA. ISSN 0018-8689.

USENIX:1982:UAS

[USE82] /usr/group, USENIX Association, Software Tools Users Group Joint Conference Proceedings: Boston, July 1982. USENIX, San Francisco, CA, USA, 1982.

VanPelt:1983:AAU

[Van83] W. V. Van Pelt. Another approach to using Writer’s Workbench programs: small class applications. In Burton and Short [BS83], pages 725–729. ISBN 0-914894-96-X. LCCN AZ105 .I56

VanOrden:1987:RIR

[Van87] G. C. Van Orden. A ROWS is a ROSE: spelling, sound and reading. Memory and Cognition, 15: 181–198, 1987. CODEN MYC-GAO. ISSN 0090-502X.

Varol:2012:HMA

[VB12] Cihan Varol and Coskun Bayrak. Hybrid matching algorithm for personal names. Journal of Data and Information Quality (JDIQ), 3(4):8:1–8:??, September 2012. CODEN ???? ISSN 1936-1955.

Volpano:1984:EIC

[VD84] D. Volpano and H. Dunsmore. Empirical investigation of COBOL features. Information Processing and Management, 20 (1/2):277–291, 1984. CODEN IP-MADK. ISSN 0306-4573 (print), 1873-5371 (electronic). From Computing Reviews: “… (6) COBOL compilers can help solve some problems by: (a) Coercing edited numeric data items in arithmetic expressions, as does FORTRAN. (b) Automatically correcting simple spelling errors with known techniques. (c) Better and more accurate diagnostics in compilers which are deficient in these areas.”.

Venezky:1976:NHE

[Ven76] Richard L. Venezky. Notes on the history of English spelling. Visible Language, 10(4):??, October 1976. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic).

Vidal:1995:FCN

[VMA95] Enrique Vidal, Andres Marzal, and Pablo Aibar. Fast computation of normalized edit distances. IEEE Transactions on Pattern Analysis and Machine Intelligence, 17(9):899–902, September 1995. CODEN ITPIDJ. ISSN 0162-8828.

Vincent:1994:DRC

[VP94] Luc M. Vincent and Theo Pavlidis, editors. Document recognition: Conference — 9– 10 February 1994, San Jose, CA, volume 2181 of Proceedings — SPIE, the International Society for Optical Engineering. SPIE Optical Engineering Press, Bellingham, WA, USA, 1994. ISBN 0-8194-1476-X. ISSN 0361-0748. LCCN TA1640.D63 1994.

Vagelatos:1994:SCS

[VTTC94] A. Vagelatos, T. Triantopoulou, C. Tsalidis, and D. Christodoulakis. A spelling correction system for Modern Greek. International Journal on Artificial Intelligence Tools [Architectures, Languages, Algorithms], 3(4):429–450, December 1994. CODEN IAITEL. ISSN 0218-2130.

Vagelatos:1995:ULS

[VTTC95] A. Vagelatos, T. Triantopoulou, C. Tsalidis, and D. Christodoulakis. Utilization of a lexicon for spelling correction in Modern Greek. Proceedings of the ACM Symposium on Applied Computing, pages 267–271, 1995.

Wagner:1974:OCR

[Wag74] R. A. Wagner. Order-n correction for regular languages. Communications of the Association for Computing Machinery, 17(5):??, May 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Witten:1983:RKN

[WCD83] Ian H. Witten, John G. Cleary, and John J. Darragh. Reactive keyboard: a new technology for text entry. Proceedings — Canadian Information Processing Society, pages 151–156, 1983. CO-DEN PCISE7.

Weiss:1982:WSE

[Wei82] E. H. Weiss. The writing system for engineers and scientists. Prentice-Hall, Inc., Englewood Cliffs, NJ 07632, USA, 1982. ISBN 0-13-971606-8. xiv + 274 pp. LCCN T11 .W44. US$16.95.

Wagner:1974:SSC

[WF74] Robert A. Wagner and Michael J. Fischer. The string-to-string correction problem. Journal of the ACM, 21(1):168–173, January 1974. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Willson:2010:ESR

[WG10] Rebekah Willson and Lisa M. Given. The effect of spelling and retrieval system familiarity on search behavior in online public access catalogs: a mixed methods study. Journal of the American Society for Information Science and Technology: JASIST, 61(12): 2461–2476, December 2010. CO-DEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Wiener:1986:EVH

[Wie86] Richard S. Wiener. Efficient virtual hash algorithm for a spelling checker. Journal of Pascal, Ada and Modula-2, 5(1):23–29, January–February 1986. CODEN JOPAD5. ISSN 0735-1232.

Williams:1982:PIS

[Wil82] M. B. Williams, editor. Pathways to the Information Society, Proceedings of the Sixth Informational Conference on Computer Communications (London, Sept. 7–10, 1982). North-Holland Publishing Co., Amsterdam, The Netherlands, 1982. ISBN 0-444-86464-4. LCCN TK5105.5 .I57 1982.

Williams:1994:NOM

[Wil94] M. E. Williams, editor. 15th National Online Meeting. Proceedings — 1994. Learned Information, Medford, NJ, USA, 1994. ISBN 0-938734-84-9. LCCN ????

Wright:1983:PTS

[WL83] P. Wright and A. Lickorish.

Proof-reading texts on screen and paper. Behav. Inf. Technol, 2(3): 227–235, July–September 1983. CODEN BEITD5. ISSN 0144-929X.

Wolff:1984:ILS

[WL84] J. G. Wolff and S. G. C.

Lawrence. Inductive learning of spelling-to-phoneme rules by data reduction. In O’Shea [O’S84], pages 506–507. ISBN 0-444-87611-1. LCCN Q334.E97 1984.

Wang:1977:SMD

[WMRB77] C. H. C. Wang, P. C. Mitchell, J. S. Rugh, and B. W. Basheer. Statistical method for detecting spelling errors in large data bases. Digest of Papers — IEEE Computer Society International Conference, ??(??):124–128, ???? 1977. CODEN DCSIDU.

Woods:1985:LPS

[Woo85] W. A. Woods. Language processing for speech understanding. In Fallside and Woods [FW85], pages 305–334. ISBN 0-13-163841-6. LCCN TK7895.S65 F351 1985. US$39.95.

Wothke:1993:MBA

[Wot93] K. Wothke. Morphologically based automatic phonetic transcription. IBM Systems Journal, 32(3):486–511, 1993. CO-DEN IBMSA7. ISSN 0018-8670. G321-5522.

Wilson:1984:DSI

[WS84] Heather Wilson and Michael J. Shaw. Designing software for the international market. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 35 (9):31–35, September 1984. CO-DEN HPJOAX. ISSN 0018-1153.

Willis:1994:SCF

[WS94] Tony Willis and Rhett Skubis.

Spell checks not foolproof. C: JET, Communication: journalism education today, 28(1):14, Fall 1994. ISSN 0198-6554.

Wang:2021:CSG

[WWD^{+}21] Yu Wang, Yuelin Wang, Kai Dang, Jie Liu, and Zhuo Liu. A comprehensive survey of grammatical error correction. ACM Transactions on Intelligent Systems and Technology (TIST), 12(5):65:1–65:51, October 2021. CODEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

URL https://dl.acm.org/doi/
10.1145/3474840.

JunWu:1994:CSU

[WWSG94] Jun Wu, Zuoying Wang, Jiasong Sun, and Jin Guo. Chinese speech understanding and spelling-word translation based on the statistics of corpus. In Anonymous [Ano94b], pages 207–210. 4 vol.

Yannakoudakis:1987:EFS

[Yan87] E. J. Yannakoudakis. Efficient file structure for specialized dictionaries and other ‘lumpy’ data. Information Processing and Management, 23(6):563–571, 1987. CODEN IPMADK. ISSN 0306-4573 (print), 1873-5371 (electronic).

Yannakoudakis:1991:GSR

[Yan91]  E. J. Yannakoudakis.  Gen-

eration of spelling rules from phonemes and their implications for large dictionary speech recognition. Speech Communication, 10(4):381–394, November 1991. CODEN SCOMDH. ISSN 0167-6393.

Yares:1995:ARL

[Yar95] Evan Yares. AutoCAD R13 looks to the future. CAE, Computer-Aided Engineering, 14(4), April 1995. CODEN CCAEDJ. ISSN 0733-3536.

Yeh:2015:CSC

[YCS15] Jui-Feng Yeh, Wen-Yi Chen, and Mao-Chuan Su. Chinese spelling checker based on an inverted index list with a rescoring mechanism. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 14(4):17:1–17:??, October 2015. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic). Special issue on Chinese spell checking.

Yoon:1991:SCI

[YEMR91] Hee Lee Yoon, Martha Evens, Joel A. Michael, and Allen A. Rovick. Spelling correction for an intelligent tutoring system. Lecture Notes in Computer Science, ??(507):77–83, 1991. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yerkey:1986:TLS

[Yer86] Neil Yerkey. Turbo Lightning: Spelling correction as you type. Electronic Library, 4(5):272–274, October 1986. CODEN ELLIDZ. ISSN 0264-0473.

Yannakoudakis:1983:ISE

[YF83a] E. J. Yannakoudakis and D. Fawthrop. An intelligent spelling error corrector. Information Processing and Management, 19(2): 101–108, ???? 1983. CODEN IP-MADK. ISSN 0306-4573 (print), 1873-5371 (electronic).

Yannakoudakis:1983:RSE

[YF83b] E. J. Yannakoudakis and D. Fawthrop. The rules of spelling errors. Information Processing and Management, 19(2):87–99 (or 87– 100??), ???? 1983. CODEN IP-MADK. ISSN 0306-4573 (print), 1873-5371 (electronic).

Yule:1994:PFR

[Yul94] Valerie Yule. Problems that face research in the design of spelling. Visible Language, 28(1):26–46, January 1994. CODEN VSLGAO. ISSN 0022-2224 (print), 2691-5529 (electronic). URL https://

s3-us-west-2.amazonaws.com/
visiblelanguage/pdf/V28N1_
1994_E.pdf.

Zamora:1980:ADC

[Zam80] Antonio Zamora. Automatic detection and correction of spelling errors in a large data base. Journal of the American Society for Information Science, 31(1):51– 57, January 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Zhao:2017:HMC

[ZCX^{+}17] Hai Zhao, Deng Cai, Yang Xin, Yuzhu Wang, and Zhongye Jia.

A hybrid model for Chinese spelling check. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 16(3):21:1–21:??, March 2017. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (electronic).

Zobel:1995:FAM

[ZD95] J. Zobel and P. Dart. Finding approximate matches in large lexicons. Software—Practice and Experience, 25(3):331–345, March 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).